### v2

# A Comprehensive Assessment Benchmark for Rigorously Evaluating Deep Learning Image Classifiers

Abstract
Reliable and robust evaluation methods are a necessary first step towards developing machine learning models that are themselves robust and reliable. Unfortunately, current evaluation protocols typically used to assess classifiers fail to comprehensively evaluate performance as they tend to rely on limited types of test data, and ignore others. For example, using the standard test data fails to evaluate the predictions made by the classifier to samples from classes it was not trained on. On the other hand, testing with data containing samples from unknown classes fails to evaluate how well the classifier can predict the labels for known classes. This article advocates bench-marking performance using a wide range of different types of data and using a single metric that can be applied to all such data types to produce a consistent evaluation of performance. Using such a benchmark it is found that current deep neural networks, including those trained with methods that are believed to produce state-of-the-art robustness, are extremely vulnerable to making mistakes on certain types of data. This means that such models will be unreliable in real-world scenarios where they may encounter data from many different domains, and that they are insecure as they can easily be fooled into making the wrong decisions. It is hoped that these results will motivate the wider adoption of more comprehensive testing methods that will, in turn, lead to the development of more robust machine learning methods in the future.
Keywords:
deep learning, neural networks, classification, generalisation, robustness, out-of-distribution, open-set, common corruptions, adversarial training, data augmentation
Code:
1 Introduction
1.1 Need for Robustness
The tremendous success of deep learning (DL) makes it easy to overlook the fact that such models are often extremely brittle and insufficiently reliable for deployment in real-world scenarios, particularly in domains where security, safety, or trust are concerns (Amodei et al.,, 2016; Heaven,, 2019; Serre,, 2019; Yuille and Liu,, 2021; Bowers et al.,, 2023; Marcus,, 2020). This lack of robustness is illustrated in Fig. 1, using the domain of image classification as an example. Essentially DL, and more generally machine learning (ML) and artificial intelligence (AI), is poor at correctly generalising to novel data/situations. However, to be dependable it is essential that such methods be able to make accurate predictions on data that was not used for training, i.e., for the trained model to generalise to novel, unseen, data.
1.2 Types of Robustness
True class=
Prediction=
IID
Generalisation
(clean accuracy)
phone
✓ phone ✓
OOD
Generalisation
(corrupt accuracy)
phone
✗ radio ✗
Adversarial Robustness
(adversarial acc.)
phone
✗ radio ✗
OOD Detection
(novel class rejection)
unknown
✗ phone ✗
OOD Detection
(unrecognisable image rejection)
unknown
✗ phone ✗
The ability to generalise can be assessed in multiple ways. Firstly, generalisation performance can be assessed using a test or validation set (see Fig. 1a). Such data is typically similar to the training data as it is drawn from the same distribution in the input feature-space. This assessment method is thus said to measure in-distribution generalisation, or generalisation to independent and identically distributed (IID) samples. Such data is routinely used to assess the performance of classifiers.
A more challenging form of generalisation is out-of-distribution (OOD) generalisation (see Fig. 1b). This can be assessed using test data that has different characteristics than the training data, but the same class labels. For example, in the domain of image categorisation it might be necessary for the classifier to cope with changes in appearance caused by images being captured under different conditions to those used to capture the training images. This can result in variations in orientation, scale, viewpoint, background, clutter, lighting, blur, jpg compression, image resolution, etc.. Furthermore, to succeed at OOD generalisation the classifier should ideally be tolerant to changes in appearance due to within-class variations, and for certain types of object (such as human bodies) non-rigid deformations. For image classification tasks, OOD generalisation can be assessed using clean samples that have been deliberately chosen because they differ from the training data (Zhang et al., 2023b, ; Hendrycks et al., 2021b, ; Hendrycks and Dietterich,, 2019; Li et al.,, 2017) or using IID test images that have been synthetically modified, or corrupted, to simulate changes in image capture conditions (Michaelis et al.,, 2019; Hendrycks et al., 2021a, ; Mu and Gilmer,, 2019; Hendrycks and Dietterich,, 2019; Jansson and Lindeberg,, 2020).
The limits placed on such generalisation may be task-dependent. For example, the image on the bottom-right of Fig. 1b shows a phone with a smashed screen. If the task were classification then this image should be identified as a phone, even if no training image showed an exemplar with a broken screen. However, if the task being performed was anomaly detection (i.e., finding defects in objects) then this exemplar would need to be identified as an anomaly. OOD generalisation is closely related to the task of domain generalisation (DG), although DG typically employs multiple training data-sets from different domains to train a method to generalise to an unseen target domain. In turn DG is closely related to domain adaptation (DA), although DA allows access to (a limited amount of) training data from the target domain to improve model transfer (Wang et al.,, 2022; Riaz and Smeaton,, 2023).
A third method of assessing generalisation performance is through the use of adversarial exemplars (see Fig. 1c). These are clean, in-distribution, samples that have been modified (or “perturbed”) in such a way as to cause a change in the predicted class even when the class of the perturbed input appears unchanged to a human observer (Akhtar and Mian,, 2018; Szegedy et al.,, 2014; Goodfellow et al.,, 2015; Kurakin et al.,, 2017; Eykholt et al.,, 2018; Biggio and Roli,, 2018). Adversarial perturbations have the appearance of low-amplitude random noise, and can be generated using a wide range of different algorithms (Liang et al.,, 2022; Ren et al.,, 2020; Xu et al.,, 2020) which attempt to find small changes in the inputs that cause large changes in the predictions of the classifier.
A fourth requirement for robust ML is the ability to reject samples from categories that were not seen during training (see Fig. 1d). Current DL models are susceptible to producing high-confidence, but incorrect, predictions to samples that are far from the training data distribution (Amodei et al.,, 2016; Hendrycks and Gimpel,, 2017). The susceptibility of a classifier to making such errors can be evaluated simply by measuring its response to samples taken from data-sets containing classes distinct from those used for training. The task of creating classifiers that are less susceptible to making such errors is called open-set recognition (OSR; Vaze et al.,, 2022; Yang et al.,, 2022) or OOD detection/rejection (Mohseni et al.,, 2020; Hendrycks and Gimpel,, 2017; Bitterwolf et al.,, 2022; Zhang and Ranganath,, 2023; Hendrycks et al.,, 2022). This task is also related to anomaly detection: as aim is to identify (and reject) novel or unexpected samples.
Finally, another situation in which an image classifier may produce high-confidence, but incorrect, predictions is when tested with exemplars that are unlike images of any object. Typically, synthetic images that occupy regions in feature-space that are far from those occupied by natural images, and hence, appear to human observers to have no similarity to the class label predicted by the classifier (see Fig. 1e). Such images can be created using stochastic processes, or like adversarial exemplars, can be algorithmically generated to deliberately fool the classifier. This latter category, known as “fooling images”, may look like noise (Hendrycks and Gimpel,, 2017), or textures (Nguyen et al.,, 2015), or images that are uniform except for a few differently coloured pixels (Kumano et al.,, 2022).
The two images at the bottom-left of Fig. 1b and d, illustrate a fundamental difficulty of accurate classification. It is necessary to have tolerance to within-class variation, and other changes in image capture conditions, so that visually dissimilar images can be allocated to the same class, while simultaneously being able to make fine-grained distinctions so that similar looking images from different classes can be distinguished (Bar,, 2004).
1.3 Terminology
OOD generalisation is concerned with classifying data where the distribution of the input (i.e., the characteristics of the images) has changed, but the distribution of output (i.e., the range of class labels associated with the images) is constant. In contrast OOD rejection is concerned with identifying data where the distribution of both the input and output has changed. The term “distribution” is therefore being used to refer to different distributions in each case. For OOD generalisation it is the distribution of the input, while for OOD rejection it is primarily the distribution of the output (the change in the input distribution is simply a side-effect of the change to new class labels). To avoid any possible confusion, OOD detection/rejection will hence-forth be referred to as “Unknown Class Rejection”. The data used to assess unknown class rejection will be subdivided into “novel” classes (images containing objects from classes unseen during training) and “unrecognisable” images (those containing no sematically meaningful information).
The term “distribution” when applied to the input is also poorly defined and, arguably, irrelevant when discussing IID/OOD generalisation. All data other than that included in the training set is OOD for the classifier (Guérin et al.,, 2022), so the distinction between IID and OOD data is qualitative rather than quantitative, and the boundary between in- and out-of distribution is necessarily arbitrary. Furthermore, the shift in distribution in image-space often has little correlation with the difficulty in correctly classifying an image. For example, two images of the same object taken from different viewpoints may have very little similarity in terms of corresponding pixels values, but the images may be easily identified as containing the same object. In contrast, it may be very difficult for the classifier to assign the same label to two images that are almost identical in image-space (e.g., a clean image and an adversarially perturbed version of that same image). Most importantly, how far an image is from the training distribution is irrelevant to the task of the classifier. Ideally, the classifier should be able to successfully recognise novel exemplars of the objects it was trained on however similar or dissimilar they are in appearance to the images used in training. In contrast an ideal classifier should reject images containing unknown objects even if such images are close to the training distribution. Hence, rather than using the terms IID and OOD, the remaining text will refer to such data as being from “known” classes. The broad class of known data will be sub-divided into different types of test data-set: “clean” data from the standard test set, “corrupt” for samples from the standard test set that have been manipulated to test for generalisation to changes in viewing conditions, and “adversarial” for samples that have been adversarially-perturbed.
The terms “generalisation” and “robustness” both refer to a classifier’s ability to produce the correct predictions. These two terms will, therefore, be used interchangeably. They will be used not only to refer a classifier’s ability to predict correct class labels for samples from known categories, but also a classifier’s ability to correctly reject samples from unknown classes.
1.4 Issues with Current Assessment Methods
While there has been a great deal of previous work carried out on assessing and improving all forms of robustness described in Section 1.2, almost all of this work has considered one form of generalisation in isolation. For example, generalisation to clean test data has been the sole pre-occupation of most research in the history of ML so far. This has resulted in methods that produce exceptional performance on test/validation data, but that are poor at generalisation to changes in appearance (Recht et al.,, 2018, 2019; Yadav and Bottou,, 2019; Geirhos et al.,, 2018; Sa-Couto and Wichert,, 2021), that are highly susceptible to adversarial attacks (Ilyas et al.,, 2019; Papernot et al.,, 2016; Akhtar and Mian,, 2018), that are prone to finding short-cuts to improve clean accuracy at the expense of learning more generally useful information (Geirhos et al.,, 2020; Amodei et al.,, 2016; Lapuschkin et al.,, 2019; Malhotra et al.,, 2020; Geirhos et al.,, 2018), and that produce high confidence predictions to samples that do not belong to any of the known categories (Amodei et al.,, 2016; Hendrycks and Gimpel,, 2017; Kumano et al.,, 2022; Nguyen et al.,, 2015). Similarly, most previous work on unknown class rejection considers this task in isolation, and does not evaluate models on clean test data let alone assess generalisation to corrupted or adversarial known classes (Yang et al.,, 2024).
Some work has concurrently considered different types of robustness. For example, most work on adversarial robustness also evaluates clean accuracy. In addition, the RobustBench benchmark also considers generalisation to corupt data (Croce et al.,, 2021), but only considers robustness to one type of data when ranking models. Other work has considered: clean accuracy and generalisation to corruptions (Vasconcelos et al.,, 2021); generalisation to input distribution shifts and unknown class rejection (Khazaie et al.,, 2022; Yang et al., 2023a, ); clean accuracy, generalisation to corruptions and unknown class rejection (Pinto et al.,, 2023; Zhang et al., 2023a, ; Averly and Chao,, 2023); adversarial robustness and unknown class rejection (Song et al.,, 2020; Feng et al.,, 2022); clean accuracy, adversarial robustness and unknown class rejection (Feng et al.,, 2022; Shao et al.,, 2020); generalisation to corruptions, adversarial robustness, and unknown class rejection (Guérin et al.,, 2023; Gwon and Yoo,, 2023); or clean accuracy, adversarial robustness and generalisation (Geirhos et al.,, 2020; Sun et al.,, 2022; Rusak et al.,, 2020; Gilmer et al.,, 2019). Hendrycks et al., 2019b considers clean accuracy, generalisation to corruption, adversarial robustness, and unknown class rejection, but each type of robustness is assessed in isolation for models that have been trained in slightly different ways. Similarly, (Mohseni et al.,, 2020) considers both clean accuracy and unknown class rejection, however, the clean accuracy of the classifier is not measured after it is modified to learn to reject unknown classes. Guille-Escuret et al., (2024) benchmarks performance using corrupt, adversarial, and unknown class images, but only considers the model’s ability to detect input distribution shifts and does not measure the accuracy with which known classes are classified.
Previous work that has considered multiple types of robustness suggests that there are trade-offs. For example, maximising clean accuracy results in poor adversarial robustness, while increasing adversarial robustness decreases clean accuracy (Zhang et al.,, 2019; Tsipras et al.,, 2019; Madry et al.,, 2018; Kannan et al.,, 2018). Similarly, many adversarial defences do not improve robustness to image corruptions and methods that improve generalisation to corrupt data typically fail to produce a strong defence against adversarial attack (Sun et al.,, 2022; Rusak et al.,, 2020; Gilmer et al.,, 2019). Improving adversarial robustness has also been shown to make a classifier less sensitive to changes in object class (Tramèr et al.,, 2020; Rauter et al.,, 2023), and less able to perform unknown class rejection (Song et al.,, 2020). While Vaze et al., (2022) show that there is generally a correlation between clean accuracy and the ability to detect unknown classes, it is not hard to imagine that methods could be developed to reject samples from unknown classes that will also result in the rejection of known samples (especially ones, such as corrupt images, that differ in appearance from the training samples) (Khazaie et al.,, 2022; Aboitiz et al.,, 2023). The converse would be a model which has been designed to be tolerant to changes in appearance in order to improve generalisation to corruptions: such a model is likely to classify more images of unknown objects with high certainty.
These trade-offs in performance on different data types raise the concern that efforts to increase robustness in one area may be resulting in a decrease in robustness in another area. However, as models are not comprehensively assessed for all different types of generalisation performance, such issues will not be noticed. Failure to fully assess the performance of models may result in huge efforts being wasted producing models that give a false sense of security by appearing to be robust only because their performance has not been tested on data for which they perform poorly.
1.5 Issues with Current Metrics
As outlined in the previous section, there is a need for a more comprehensive assessment of robustness, using multiple different types of test data-set as illustrated in Fig. 1. However, metrics that are most commonly used to assess each individual form of robustness are inadequate for producing more comprehensive and rigorous assessment.
Performance on known classes (clean, corrupt, and adversarial samples) is typically evaluated by calculating the classification accuracy (or equivalently the percentage error), see Fig. 2 and Table 1, although many other metrics can also be used (Guo et al.,, 2023). However, if models are to be allowed to reject samples, to deal with those from unknown classes, then the performance on known samples should also be assessed taking into account that some samples may be rejected. On the other hand, models designed to perform unknown class rejection (Kirchheim et al.,, 2022; Chen et al.,, 2023; Cheng et al.,, 2023; Xu-Darme et al.,, 2023; Yang et al.,, 2022; Yang et al., 2023b, ; Lee et al.,, 2022), and adversarial defences that reject perturbed samples (Stutz et al.,, 2020; Ren et al.,, 2022; Lee et al.,, 2022; Chen et al.,, 2022), are typically evaluated purely in terms of how accurately samples are rejected and accepted. In other words, the task is considered to be that of binary classification (see Fig. 2 and Table 1). Hence, standard metrics for binary classification can be used, such as the False Positive Rate when the true positive rate is X% (FPR@X%), which assesses performance for a single rejection threshold, or Area Under the Receiver Operating Characteristic curve (AUROC) or Area Under the Precision-Recall curve (AUPR) both of which assess performance over varying rejection thresholds. An issue with FPR@X% is that separate thresholds are used for distinguishing each unknown data-set from the clean test data. Worse still, some rejection methods have hyper-parameters that are tuned separately on each unknown data-set (Zhang et al., 2023b, ). This is unrealistic, as in the real-world a single rejection criteria with fixed hyper-parameters would need to work to reject any unknown data, regardless of its origin (Shafaei et al.,, 2019; Feng et al.,, 2022). An issue with all standard unknown class rejection metrics is that the ability of the model to accurately classify the accepted samples is not evaluated. In some publications the clean accuracy of the model is evaluated, but this is done without any samples being rejected, and hence, does not reflect the performance that would be expected in the real-world where the same rejection criteria would be applied to all samples.
| predicted class | |||||
| 1 | 2 | … | c | ||
| true class | 1 | ✓ | ✗ | ✗ | |
| 2 | ✗ | ✓ | ✗ | ||
| c | ✗ | ✗ | ✓ | ||
| predicted | |||
| accept | reject | ||
| true | known | ||
| unknown | |||
| predicted class | |||||
| correct | wrong | ||||
| true | known | predicted | accept | ||
| reject | |||||
| unknown | accept | N/A | |||
| reject | N/A | ||||
Some metrics do attempt to evaluate classification performance in the presence of rejection. For example, in the context of evaluating accuracy on adversarial data, confidence-thresholded robust error (Stutz et al.,, 2020) measures the classification error for test examples that are not rejected, when the rejection threshold has been set so that 1% of correctly classified clean examples can be rejected. Using the metric robust risk with detection (Tramèr,, 2021) a classifier is accurate if it correctly classifies the clean samples, and either rejects or correctly classifies perturbed samples. However, with these metrics a high score can be obtained by a classifier that rejects a large proportion of the perturbed samples, even though these samples come from a known class and may be only perturbed imperceptibly (Chen et al.,, 2022). To address this issue, robustness with rejection (Chen et al.,, 2022) defines a variable perturbation threshold that can vary from zero to the adversarial attack budget. For test samples with a perturbation less than or equal to the threshold, both incorrect classification and rejection is counted as an error. For test samples with a perturbation greater than the threshold, incorrect classification is considered an error, but rejection is not counted as an error. When the threshold is greater or equal to the attack perturbation budget, robustness with rejection is equivalent to standard robust error. When the threshold is zero this metric is equivalent to robust risk with detection as proposed by Tramèr, (2021).
An issue with all the methods discussed in the previous paragraph, is that they define which samples can be rejected in terms of an arbitrary measure of distribution shift in the input space (Guérin et al.,, 2023). This means that these methods would not allow for rejection of non-adversarial samples. For example, in the situation where the sample is clean but ambiguous (e.g., in an optical character recognition task a letter written in terrible handwriting) rejection would be counted as an error even though the distribution shift for such a sample might be far higher than that of any adversarial sample. These metrics would also be incapable of assessing performance on data-sets containing unknown samples. In this situation the correct action of the classifier is to reject all samples, however, for confidence-thresholded robust error (Stutz et al.,, 2020) a score is only calculated for accepted samples, so no score could be calculated for a classifier correctly rejecting all unknown samples. Both robustness with rejection (Chen et al.,, 2022) and robust risk with detection (Tramèr,, 2021) would suffer from the same issue, assuming unknown samples were considered to be sufficiently perturbed to be out-of-distribution. Alternatively, if clean samples from unknown classes were considered to be unperturbed, these metrics would give a score of zero to a classifier that correctly rejected all samples, as any rejection of clean sample is counted as an error for these metrics.
It would be more reasonable to decide whether or not rejection or acceptance of the sample was the correct choice depending on whether or not that sample would have been mis-classified or correctly classified (Pang et al.,, 2022; Averly and Chao,, 2023), as the goal of sample rejection is to protect the classifier from making predictions that are unreliable, not to identify samples that differ from the training data (Guérin et al.,, 2023; Zhu et al.,, 2024). Taking this approach, Zhu et al., (2024) propose a metric that they apply to evaluate the performance of classifiers using both corrupt samples and unknown samples from novel classes. As shown in Fig. 2 and Table 1, unlike standard metrics, this approach considers both whether or not a sample is rejected, and whether or not the classification label predicted by the classifier is correct. In this scheme, a sample is considered a true-positive () if it is not rejected and assigned the correct class label. A sample is a false-positive () if it is accepted but incorrectly classified. A false-negative () is a sample that is rejected although it would have been assigned the correct class label. Finally, true-negatives () are samples that are rejected and that would have been classified incorrectly. The proposed metric is the detection error rate (DER), which counts the number of false positives and false-negatives as a proportion of all test samples. Zhu et al., (2024) apply their metric using two rejection thresholds, that are set such that either 95% or 99% of correctly classified clean examples are accepted. These different values for the rejection threshold define the trade-off between accuracy on the clean data (perfect performance on the clean data could only be achieved if 100% of correctly classified clean examples were accepted) and correct rejection of unknown samples (good performance on unknown class rejection is more likely if more clean samples are also rejected).
2 Methods
2.1 Proposed Assessment Method
This article proposes to extend and modify the evaluation method proposed by Zhu et al., (2024), described in the preceding section, in the following ways:
1.
Measuring performance in terms of accuracy rather than error. The metric defined by Zhu et al., (2024), DER, is a measure of the proportion of errors (expressed as a percentage) made by the classifier. Here, the percentage of correctly processed samples is reported. This simple change means that a higher score correlates with improved performance which is more intuitive, and is more consistent with previously used metrics (such a AUROC and those used in RobustBench (Croce et al.,, 2021)) where this is also the case. Hence, rather than using detection error rate (DER), detection accuracy rate (DAR) is reported.
2.
Summarising results by averaging over task, not data-set. Zhu et al., (2024) report results for each data-set separately. However, using many separate metrics makes it difficult to evaluate the overall robustness of any model especially when there are trade-offs in performance of different data-sets (meaning that no method of improving robustness is likely to have strong performance across all data-sets). It is therefore convenient to have a single metric to summarise performance. For this purpose, Zhu et al., (2024) average the performance across each data-set. However, there is a risk that a more extensive evaluation of robustness using more data-sets might result in this average becoming biased towards one type of robustness. For example, as there are potentially a very large number of data-sets containing images from unknown classes, including more such data-sets would result in the average performance favouring methods better able to reject unknown classes. Alternatively, calculating a single value of the metric using all data from all data-sets would be biased towards favouring methods better able to deal with data-sets containing the most samples. To better balance the importance of robustness to each type of data, here a single summary metric is obtained by calculating the metrics separately for each type of data, and then averaging across data type aaaAs the sub-sets of data used for each type of data (see Section 2.2) contain similar numbers of exemplars, the DAR for each type of data is calculated for all the data of that type. However, if the metric was to be used in future with less well-balanced sub-data sets, it would be sensible to calculate the DAR speartely for each sub-set and then average over those results to produce the estimate of DAR for that data type..
3.
Extending the evaluation to consider other forms of generalisation. Zhu et al., (2024) evaluate performance using several data-sets including clean and corrupt data and data from unknown classes. In the current work, unknown class rejection is evaluated not only using images from novel classes, as in (Zhu et al.,, 2024), but also using unrecognisable images. Furthermore, performance is also evaluated using adversarial samples. In other words, here performance is evaluated using the five types of data illustrated in Fig. 1, whereas Zhu et al., (2024) use only three.
2.2 Choice of Test Data
The current work tested the performance of Deep Neural Networks (DNNs) trained on the following data-sets: CIFAR10, CIFAR100, TinyImageNet, and MNIST. In each case training was performed with the standard training sub-set of data. The robustness of each trained network was tested using the five categories of test data illustrated in Fig. 1. Within each category the following data was used.
- Clean.
The standard test-set provided with each data-set.
- Corrupt.
Popular common corruptions data-sets were used: CIFAR10-C, CIFAR100-C, TinyImageNet-C and MNIST-C (Hendrycks and Dietterich,, 2019; Mu and Gilmer,, 2019). The first three data-sets contain 18 different corruptions including different types of noise, blurring, synthetic weather conditions, and digital corruptions. MNIST-C contains 15 different corruptions including different types of noise, blurring, geometric transformations, and superimposed patterns. As typical in the literature, performance was evaluated using all the corruptions at all degrees of intensity.
- Adversarial.
Samples were produced using AutoAttack (AA; Croce and Hein,, 2020), a state-of-the-art ensemble attack method that employs both gradient-based (white-box) and gradient-free (black-box) attacks. AA was implemented using the torchattacks PyTorch library (Kim,, 2021). Two sets of adversarial samples where created. Each set was created by perturbing the entire standard (clean) test-set, but with a different method of constraining the magnitude of the perturbation. Specifically, AA was used to apply both and -norm constrained attacks. The perturbation budget () used for each attack was the standard value used in the previous literature for each data-set. Specifically, was set to and 0.5 for and -norm attacks, respectively, against networks trained on the CIFAR10, CIFAR100, and TinyImageNet data-sets, and was set to 0.3 for -norm and to 2 for -norm attacks on MNIST trained networks.
- Novel Class.
In each case, three data-sets containing objects of unknown classes were chosen. Specifically, the CIFAR10 and CIFAR100 trained networks were tested using novel classes from the test-sets of the Textures (Cimpoi et al.,, 2014) and SVHN (Netzer et al.,, 2011) data-sets and the CIFAR data-set that was not used for training (i.e., a CIFAR10 trained network was evaluated on its ability to reject samples from CIFAR100, and vice versa). For TinyImageNet trained networks novel class data came from the Textures test set (Cimpoi et al.,, 2014), the iNaturalist 2021 validation set (Van Horn et al.,, 2018), and the ImageNet-O data-set (Hendrycks et al., 2021c, ). The MNIST trained networks were tested using unknown, novel, classes from the test-sets of the Omniglot (Lake et al.,, 2015), FashionMNIST (Xiao et al.,, 2017) and KMNIST (Clanuwat et al.,, 2018) data-sets.
- Unrecognisable.
Data to test unrecognisable image rejection was the same for all networks, and consisted of synthetic images randomly generated using four different methods. First, images containing random blobs, as used in (Hendrycks et al., 2019a, ). Secondly, images in which each pixel intensity value was independently and randomly selected from a uniform distribution. Thirdly, the images of the standard (clean) test set after a random permutation of all pixels. Fourthly, the images of the clean test set after randomising the phase, in the Fourier domain, of each image. Each of these four data-sets contained the same number of samples as the standard test-set associated with the training data.
In principal, the amount of test data that could be used is unlimited (as there are countless images containing objects that any one classifier has not been trained to recognise, and due to the stochastic nature of the methods used to create corrupt, adversarial, and unrecognisable images). Hence, the above selection of test data necessarily represents a small, arbitrary, sub-set of all the possible data that could be used. In making this selection the aim was to define a reasonably wide variety of data while ensuring that evaluation could be performed in a reasonable time using reasonable resources. The results in Section 3 suggest that the chosen data is sufficiently challenging that modern DNNs are far from ceiling performance. However, as classification methods improve in the future, the data used for evaluation could be extended to allow a more thorough and challenging assessment in two ways: adding more data in each category, or by adding categories.
Expanding the range of data in each of the existing categories could be done by: using new sets of clean data (such as CIFAR10.1 (Recht et al.,, 2018), or ImageNet-A (Hendrycks et al., 2021b, )); using different corruptions (e.g., different forms of noise or image transformation), higher magnitude corruptions (e.g., a wider range of scales such as used in the MNIST Large Scale data set (Jansson and Lindeberg,, 2020)), the different types of corruption used in the OOD-CV-v2 benchmark (Zhao et al.,, 2024), or objects shown against different backgrounds (as is used in the NICO challenge (Zhang et al., 2023b, )); using any of numerous alternative algorithms for generating adversarial images (Liang et al.,, 2022; Ren et al.,, 2020; Xu et al.,, 2020), or different methods of constraining the perturbations, or higher perturbation budgets; using the additional novel class data-sets proposed in the OpenOOD benchmark Yang et al., (2022); using fooling images (Kumano et al.,, 2022; Nguyen et al.,, 2015) that are unrecognisable images which have been optimised to be difficult to reject. Adversarial attack-type techniques could also be used not only to perturb clean test samples, as is currently the case, but samples from other data types to make these types of data more challenging to correctly process. New categories of test data that could be used to extend and increase the challenge of the proposed benchmark include: known classes presented in the form of drawings, art, or cartoons (such as in ImageNet-R (Hendrycks and Dietterich,, 2019) or PACS (Li et al.,, 2017)); or images where texture has been replaced by that from another image (Geirhos et al.,, 2019).
2.3 Rejection Criteria
In order for classifiers to be evaluated using the proposed method, it is necessary to define a criteria that will be used to determine if a sample is accepted for classification, or rejected. Many different methods for unknown class rejection have been proposed (Vaze et al.,, 2022; Pang et al.,, 2022; Chen et al.,, 2022; Hendrycks and Gimpel,, 2017; Tajwar et al.,, 2021). This article will primarily use maximum softmax probability (MSP; Hendrycks and Gimpel,, 2017), which defines the confidence that a sample is of a known class as the maximum response of the network output after application of the softmax activation function. MSP is used as this is a standard, baseline method, used frequently in the literature on unknown class rejection. Furthermore, a comparison of different rejection methods using the same metric on which the proposed benchmark is based found MSP to be one of the best methods (Zhu et al.,, 2024).
Previous work has shown that relative performance of different unknown class rejection methods vary depending on a number of factors (Szyc et al.,, 2023) including the data-sets Tajwar et al., (2021) and models (Zhu et al.,, 2024) used. Hence, to confirm that the results obtained in this study are not unique to MSP results for three other post-hoc methods of unknown class rejection (those that rely on the responses produced by the network) are also reportedbbbOnly post-hoc methods are considered, as (1) they can be applied to the same network, allowing for a fair comparison and avoiding the need to re-train the network to assess different methods, (2) because post-hoc methods have performance comparable to, or better than, other methods (Yang et al.,, 2022). (1) Maximum logit score (MLS; Vaze et al.,, 2022), which defines the confidence that a sample is of a known class as the maximum response of the network output before any activation function is applied. This was shown to be the best method of unknown class rejection in a recent comparison of state-of-the-art methods (Vojir et al.,, 2023). (2) Energy score (Liu et al.,, 2020), which defines confidence as the negative logarithm of the denominator of the softmax activation function applied the network output layer. (3) GEN (Liu et al.,, 2023) a very recently proposed method that claims to be superior to all three other methods considered here.
Consistent with the method proposed by Zhu et al., (2024), a rejection threshold is set such that a fixed percentage correctly classified clean examples are accepted. In other words, such that the confidence associated with those samples (calculated using either MSP, MLS, Energy score or GEN depending on the experiment being performed), is greater than or equal to the threshold. All samples, from any data-set, with confidence lower than the threshold are rejected, while all others are accepted and the class label predicted.
2.4 Choice of Models to Assess
The aim of the current work is to determine how well existing DNNs perform when evaluated using the more comprehensive assessment methods advocated above. However, it is clearly impossible to attempt to evaluate the performance of every network architecture or to evaluate the huge variety of different techniques that has been proposed for improving adversarial robustness, generalisation to common corruptions, or for unknown class rejection.
ResNet18 (He et al.,, 2016) was chosen as the principal DNN architecture to test as this is very commonly used in the literature. This model has just over 11M trainable parameters. Additional experiments were also performed using WideResNet34-20 (Zagoruyko and Komodakis,, 2016) which has just under 193M parameters. For the MNIST data-set smaller networks were used. Firstly, a full-connected multi-layer perceptron network, containing 200 neurons in each of 3 hidden layers (denoted as MLP200-200-200). This network used the ReLU activation function and had 239,410 trainable parameters. Secondly, a fully-convolutional neural network (ConvNet) containing 5 layers, with 32 3x3 convolutional filters in each hidden layer, and average pooling of stride 2 applied after the first and second layers. The ReLU activation function was applied to each layer (except the final output). This network is denoted as ConvNet5-3-32 where the numbers indicate the depth, kernel-size, and width respectively. This network contained 30,954 trainable parameters.
A common approach to improving performance on all the different types of test-data described in Section 1.2 is to use data augmentation techniques: i.e., to expand the training data-set by creating new samples that are synthetically manipulated versions of samples from the original training data-set. In the results section (Section 3) a number of such techniques are evaluated. The selection of the particular methods to test was made so as to include standard baselines and state-of-the-art methods for producing robust models for all the different types of test-data described in Section 1.2.
- baseline
Improved performance on the standard test data is often achieved by applying simple geometric data augmentations to the training data. Here 4-pixel random cropping and random horizontal flipping where applied to the CIFAR10, CIFAR100 and TinyImageNet training data, while no augmentations were applied to the MNIST data. These standard augmentations were used in combination with the additional augmentation methods listed below. Given that these baseline augmentation methods are very commonly used when training models, the results show how standardly trained networks cope under comprehensive evaluation. The results for this method of training also provides a baseline against which to judge the effectiveness of other methods that have been proposed to improve generalisation performance.
- noise
Augmenting training images with random noise has been shown to reduce over-fitting, and hence, improve generalisation (Sietsma and Dow,, 1991; Holmstrom and Koistinen,, 1992; Goodfellow et al.,, 2016), especially to common image corruptions (Rusak et al.,, 2020; Gilmer et al.,, 2019). Various methods of adding noise to the training data have been proposed. Rusak et al., (2020) augmented half the training data with Gaussian noise, and found that this technique produced performance on corrupt test data that was competitive with the more complex augmentation method, AugMix, and that for the MNIST data-set specifically was just as effective as their proposed method of Adversarial Noise Training. Lopes et al., (2019) advocated adding Gaussian noise to only a randomly chosen patch of the image, which they claimed produced better robustness to image corruptions on CIFAR10 data than applying random noise to the whole image. Gilmer et al., (2019) used an alternative method for corrupting training images with Gaussian noise in which the standard deviation of the noise applied to each image was chosen with equal probability from the range . The method implemented here is a generalisation for the method proposed by Gilmer et al., (2019). Specifically, for each image the standard deviation of the noise was chosen from a uniform distribution with range . Then each pixel intensity value was independently modified by the addition of a value chosen at random from a Gaussian distribution with mean of zero and the chosen standard deviation. The original images had pixel intensity values in the range [0,1], and after the application of the noise pixel values were clipped at 0 and 1 to remain in this range. Appropriate values of and were found via a grid search using small ConvNets. Using this method and were chosen for use with CIFAR10, and the same hyper-parameters were also used for CIFAR100 and TinyImageNet. For MNIST was used.
- AT
Multi-step adversarial training with Projected Gradient Descent (PGD; Madry et al.,, 2018) is a standard, and highly effective, defence against adversarial attack. As a result, this method has become a standard benchmark against which all other methods of adversarial defence are judged. This method was, therefore, also selected as a data augmentation technique to be tested. It was implemented using the torchattacks PyTorch library (Kim,, 2021). To perform adversarial training, clean images from the current batch of training examples were used to generate non-targeted adversarial examples, that then replaced the original training images (as in Madry et al.,, 2018). Following standard protocols, the maximum allowed perturbation was constrained by the -norm to be less than the perturbation budget (), and the training-time adversarial images were generated using 10 steps of PGD (denoted as PGD for brevity). Standard values used in the previous literature were used for . Specifically, was set to for performing AT with the CIFAR10, CIFAR100 and TinyImageNet training data, and was used for MNIST.
- pixmix
This is a training data augmentation pipeline that repeatedly mixes (either additively or multiplicatively) an augmented image with an augmented clean image or a fractal image (Hendrycks et al.,, 2022). It is claimed that this augmentation method can improve generalisation to both clean and corrupt data, improve adversarial robustness, and improve unknown class rejection. This method of data augmentation was therefore also tested using the proposed, comprehensive, evaluation method. It was implemented using the official code released by the original authors.
- regmixup
In addition, evaluations were also performed on current state-of-the-art robust pre-trained models available from the RobustBench model zooccchttps://github.com/RobustBench/robustbench (Croce et al.,, 2021).
2.5 Experimental Setup
Training was performed using a batch size of 128. For the CIFAR10, CIFAR100 and TinyImageNet data-sets, training was performed for 110 epochs using the SGD optimiser with weight decay of 5e-4 and an initial learning rate of 0.1. The learning rate was step-wise decayed by a factor of 10 at the start of epochs 100 and 105 for the CIFAR data-sets, and at epochs 45 and 89 for TinyImageNet. The set-up for the CIFAR data-sets was based on that recommended for AT by Pang et al., (2021). Hence, these hyper-parameters may not be optimal for training with the other data augmentations used. For the MNIST data, 20 epochs of training using the Adam optimiser and a constant learning rate of 0.001 was used.
For each experimental condition (combination of data-set, network architecture, and training data augmentation method) the network was trained and evaluated three times, each time with a different random weight initialisation and random presentation order of training samples. The reported result for each condition is the mean accuracy over these three trials. The standard-deviation in the mean performance measured using the proposed metric is shown in the tables of results and generally shows a high degree of consistency across trials. Exceptions to the above, where only a single trial was performed, were the experiments with WideResNet34-20 and the experiments performed using a publicly available checkpoint created by other researchers.
Open-source code written in Python using the Pytorch library (Paszke et al.,, 2019), which performs all the experiments described in this article, is available for download from: https://codeberg.org/mwspratling/RobustnessEvaluation.
3 Results
3.1 Comparison with Standard Metrics
| Training | Clean | Corrupt | AA | AA | Textures | SVHN | CIFAR100 | Phase | Scramble | Blobs | Uniform |
| Method | Accuracy (%) | AUROC (%) | |||||||||
| baseline | |||||||||||
| noise | |||||||||||
[…]
| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
| baseline | .77) | |||||
| noise | .99) | |||||
[…]
| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
| baseline | .30) | |||||
| noise | .30) | |||||
[…]
| Training Method | Clean | Corrupt | AA | AA | Textures | SVHN | CIFAR10 | Phase | Scramble | Blobs | Uniform |
| Accuracy (%) | AUROC (%) | ||||||||||
[…]
| noise | |||||||||||
[…]
| regmixup | |||||||||||
| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
[…]
| noise | .45) | |||||
[…]
| regmixup | .28) |
| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
[…]
| noise | .09) | |||||
[…]
| regmixup | .97) |
Table 2 shows results for ResNet18 networks trained on CIFAR10 using different training-time data augmentation methods that have been proposed to improve performance (as discuss in Section 2.4). Table 3 shows equivalent results for CIFAR100 trained networks. Evaluation of the performance of these networks using standard metrics is shown in Figs. 3(a) and 3(d). It can be seen that, as expected, AT with PGD produces the best defence against adversarial attacks with AA. It can also be seen that, compared to a network trained with the baseline augmentations, AT causes a reduction in the accuracy with which clean data is categorised and a reduction in performance on unknown class rejection. Also not unexpectedly, pixmix produces the best performance on unknown class rejection for all the data-sets (except CIFAR10 for the CIFAR100 trained network), and produces the best performance on the corrupt data.
These findings are also captured by the new evaluation metric, results for which are shown in Figs. 3(b) and 3(e) when the rejection threshold is set such that 99% of correctly classified clean examples are accepted. The results summarising performance on each type of data show that AT produces the highest accuracy for adversarial samples, and pixmix produces the highest accuracy on corrupt data and the two types of unknown class data (novel and unrecognisable). The trade-offs in performance over different types of data can also be seen. The average performance over all data types (shown in the last columns of each table) suggests that pixmix produces the best performance overall.
The results for all methods on unknown class data is far worse than the standard metrics would imply. This is because the same rejection threshold has been used for all data types, rather than being tuned to optimise performance on unknown class rejection. This illustrates that current metrics for unknown class rejection grossly over-estimate likely real-world performance (as discussed in Section 1.5). In contrast, the performance of all methods on known class data (clean, corrupt and adversarial) is higher than would be expected from the standard metrics. This is because networks can reject samples for which prediction confidence is low which include many samples for which the predicted class is wrong.
Note that these results generally show that current methods of training deep networks produce classifiers that lack robustness. For most of the methods evaluated (which includes methods that are designed to produce state-of-the-art robustness) the resulting network can only make the correct decision on around 50% of the samples for CIFAR10, which is a relatively simple data-set. Even pixmix, that performs best overall, only makes accurate decisions for just over 71% of the test data. Furthermore, it might reasonably be argued that a model is only as robust as the worst results would suggest, as a malicious actor can select the type of data used to attack a system, and in the real world a model might naturally encounter all types of data and would need to deal correctly with all. For each method of training investigated here there is at least one type of test data on which the model performs pitifully. For example, AT can only correctly reject 2.2% of unrecognisable images: all the others are given incorrect class labels with a high confidence. Massive amounts of time and resources have gone into developing techniques to enhance adversarial robustness. The current results suggest that this effort may have been wasted, as an attacker can easily and effectively induce a classifier to generate the wrong predictions by presenting it with easily generated synthetic images. The results in Section 3.5 show that the most recent state-of-the-art methods of adversarial robustness, while having better performance on unknown class rejection, are still far from robust to such data.
Figs. 3(c) and 3(f) show results for the proposed evaluation method when the rejection threshold is set such that 95% of correctly classified clean examples are accepted (an increased rejection threshold compared to previously). For the CIFAR10 trained networks, rejecting more of the clean samples results in a reduction in performance on the clean data as some samples that were correctly classified have been rejected. In contrast, the results for CIFAR100 show that performance on the clean data improves with the increased rejection threshold, due to the higher proportion of clean samples that are mis-classified by the CIFAR100 trained networks. For both data-sets, rejecting more samples for which confidence in the prediction is low results in a large improvement in performance on the adversarial, novel and unrecognisable data types for all training methods.
The results produced with both thresholds are consistent showing similar trade-offs in performance on the different data types for each of the tested methods, and producing similar rankings of different training methods, although the differences between methods is somewhat reduced when the threshold is set so as to accept 95% of correctly classified clean samples. To avoid redundancy, and for brevity, only results for one threshold are reported subsequently. The 95% threshold was chosen in order to present the most favourable performance for the tested DNNs.
3.2 Different Unknown Class Rejection Methods
| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std | |
| MLS | baseline | .33) | |||||
| noise | .59) | ||||||
[…]
| regmixup | .20) | ||||||
[…]
| noise | .65) | ||||||
[…]
| regmixup | .15) |
| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std | |
[…]
| noise | .95) | ||||||
[…]
| regmixup | .34) | ||||||
[…]
| noise | .16) | ||||||
[…]
| regmixup | .12) | ||||||
[…]
| noise | .05) | ||||||
[…]
| regmixup | .76) |
The same networks evaluated in the previous section were re-evaluated using MLS, Energy score and GEN as the criteria for rejection (see Section 2.3). The results for the networks trained on CIFAR10 are shown in Table 4, and for CIFAR100 in Table 5. Comparing the results in Table 4 with the results obtained using MSP shown in Fig. 3(c) it can be seen that while individual accuracy values can vary, a similar overall performance is produced using each method of rejection, and as a result a similar ranking of the overall performance of each training method is obtained. The same can be said for the CIFAR100 results shown in Table 5 and Fig. 3(f). The results obtained, therefore, do not seem to be specific to the particular method of rejection used, and to avoid redundancy, and for brevity, only results for MSP are reported subsequently. This choice was made as the worse performance on any data type is better for MSP than for any of the other three methods, and hence, MSP seems to have a slight advantage in improving robustness.
| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean |
[…]
| regmixup |
| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean |
| baseline | ||||||
| noise | ||||||
[…]
| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
| baseline | .00) | |||||
| noise | .01) | |||||
[…]
| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
| baseline | .12) | |||||
[…]
| AT | .21) | |||||
[…]
| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
| baseline | .30) | |||||
[…]
3.3 Effects of Increasing Network Capacity
An increase in network capacity is typically accompanied by an increase in performance. Hence, in an attempt to improve on the results presented in Section 3.1 these experiments were repeated using the WideResNet34-20 architecture, that has more than 17 times the number of trainable parameters compared with the ResNet18 architecture used previously. The results for CIFAR10 are shown in Table 6. Comparing these results to those produced under equivalent conditions with ResNet18 shown in Fig. 3(c), it can be seen that, surprisingly, results for regmixup are worse for the network with higher capacity, due to a poorer ability to reject samples from unknown classes. In contrast, the overall performance improves when baseline, noise, AT, and pixmix are used as the training data augmentation methods. The improvement for pixmix is only very modest, while that of AT is the largest. The large improvement in the overall performace for AT is accompanied by a large increase in robustness to unrecognisable data. However, the results on WideResNet34-20 show that for this training method, as for all the others, there is at least one type of test data on which the model performs poorly.
Comparing the results for CIFAR100 in Table 7 and Fig. 3(f) it can be seen the the performance produced by WideResNet34-20 is superior to that produced by ResNet18 for all the different training methods, and the largest increase in performance is for regmixup. However, as for the CIFAR10 results, even with a large capacity network the robustness of the networks produced by each training method is still poor and there is still at least one type of test data for which performance is very poor.
3.4 Results for Other Data-sets
The determine how well the preceding results generalise, experiments were repeated using a more challenging data-set (TinyImageNet) and an easier one (MNIST). For MNIST different network architectures were also used. The results for TinyImageNet (Table 8) are very similar to the equivalent results produced for CIFAR100 (Fig. 3(f)), but with a small drop in overall performance for each training method. In contrast, the results for MNIST (Tables 9 and 10) differ from the results for the other data-sets and differ in details between the two architectures used. Specifically, the best overall robustness is obtained for the MNIST trained networks, as this is the least challenging data-set. However, surprisingly the best performance is not produced by pixmix, as for the other data-sets, but by training with noise.
3.5 Benchmarking Leading Methods from RobustBench
| Model | Robustness | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean |
[…]
| Bartoldson et al., (2024) | adv. () | ||||||
[…]
| Model | Robustness | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean |
[…]
| Cui et al., (2024) | adv. () |
RobustBench (Croce et al.,, 2021) is an existing benchmark for assessing the robustness of DNNs. Unlike the proposed benchmark, RobustBench evaluates performance of a model using only a single type of data: either adversarial (produced using either or constrained AA) or common corruptions. The two best performing models from each RobustBench category (for which pretrained models were available from the RobustBench repository on November 15th 2024) were evaluated using the proposed benchmarkdddAA crashed when evaluating the second best method for CIFAR100 adversarial robustness (Bai et al.,, 2024) due to the model producing NaNs. This was the case when testing using AA implemented using the torchattacks library (as has been used to adversarially attack all other methods assessed in this paper), and when using the autoattack library’s version of AA (the implementation used by RobustBench). Results for the first and third best methods are therefore presented for CIFAR100 adversarial robustness.. The results are shown in Tables 11 and 12 for the CIFAR10 and CIFAR100 data-sets, respectively. Note that these methods use the following model architectures: WideResNet18-2 (268M parameters) (Diffenderfer et al.,, 2021); ResNet18 (11M parameters) (Kireev et al.,, 2022; Modas et al.,, 2022); WideResNet94-16 (366M parameters) (Bartoldson et al.,, 2024); MeanSparse RaWideResNet70-16 (267M parameters) (Amini et al.,, 2024); WideResNet70-16 (267M parameters) (Wang et al.,, 2023; Rebuffi et al.,, 2021); WideResNet28-10 (36M parameters) (Cui et al.,, 2024).
It can be seen that the state-of-the-art RobustBench methods suffer from the same issues observed in the previous experiments. For the models trained to generalise to common corruptions the performance against adversarial attacks is poor. While the models trained to be robust to adversarial attacks are poor at rejecting unrecognisable images. Hence, overall, these models lack the robustness that conventional methods of evaluation suggest they have. On the contrary, the proposed comprehensive evaluation shows that methods of improving performance against adversarial attack, for example, fail to provide security as an attacker can use cheaply and easily generated random images to fool the classifier with a high chance of success.
4 Discussion
The aim of testing a classifier is to ensure that the decision boundaries that it defines have been positioned correctly. However, as illustrated in Fig. 3, each of the types of test data illustrated in Fig. 1, on its own, is incapable of comprehensively evaluating the accuracy with which the classifier has positioned the decision boundaries for each class. For simplicity, the figure shows only two known classes, but the same observations hold true for multi-class classification problems. The clean and corrupt test data-set (due to their finite size) can not test that the classifier responds correctly to every possible exemplar of each known class. Furthermore, they can not ensure that the classifier fails to assign class labels to samples that are outside the decision boundaries of any known class. For the example shown in the figure, it would be possible for a classifier to achieve perfect performance on the clean and corrupt data by defining a straight decision boundary roughly dividing the feature-space into two halves, and entirely ignore all other boundaries around the regions occupied by the known classes. The adversarial samples evaluate in fine detail the placement of the decision boundaries between known classes, but fails to ensure that other boundaries are defined appropriately. The unknown test data (due to its finite size) can not test that the classifier responds correctly to every possible novel or unrecognisable sample. Furthermore, this type of test data can not ensure that the classifier assigns correct class labels to samples from known classes. This article advocates testing classifiers with all these types of data in order to make steps towards a more complete and comprehensive evaluation of a classifier’s positioning of the decision boundaries, and hence, its ability to make accurate predictions on unseen data.
Despite these limitations, current practice is to test classifiers using only one (typically the clean test data), or two (clean and adversarial, or clean and novel) types of test data. While others have previously advocated using more types of test data (see Section 1.4) they have proposed using different metrics for different data-sets (Section 1.5) , which fails to adequately test true performance. This article promotes an existing metric (Zhu et al.,, 2024) that is capable of being used with different types of data to provide a consistent and comprehensive evaluation of performance.
Using the proposed benchmarkeeeNote that the proposed evaluation benchmark is likely to be imperfect, and is likely to be biased: like all assessment methods it will favour certain models over others, those that are good at processing the specific data-sets that has been chosen for performing the evaluation, but may be poorer at dealing with other data. However, the proposed benchmark is less biased than those that make no attempt to evaluate robustness to certain types of data, and (as discussed in Section 2.2) the data sets used for assessment in the proposed benchmark can easily be extended. it is found that existing methods of training DNNs, including methods that are claimed to produce state-of-the-art robustness, fail to perform as well as previous, less comprehensive, assessments would suggest. Furthermore, it appears that there are trade-offs that mean that training methods that have good performance on one type of data (i.e., improve the placement of the decision boundary in one region of feature-space) suffer from poor performance on other data (i.e., place the decision boundary less accurately elsewhere in feature-space). This means that all methods of training DNNs that have been tested here are highly susceptible to malicious attacks and, by appropriate choice of data type, can be easily fooled into producing the wrong predictions. Improved performance might be possible by combining current methods: e.g., using one model to perform unknown class rejection, and an adversarially robust model to classifying the unrejected samples. However, appropriate evaluation protocols, such as the one advocated here, will still be needed to test such ensemble methods.
The lack of robustness in current DNNs is an inconvenient truth that has been, and is likely to continue to be, ignored by many as it gets in the way of claiming success on a particular task and publishing results that can be said to be state-of-the-art on the very specific data-set that has been used for evaluation. Changing this culture is likely to be only brought about by reviewers and editors insisting on more comprehensive evaluation of machine learning methods. Just as it is unacceptable to test a model on the training data, so it should become unacceptable to test a model on a single, or a couple of, types of test data.
Without such a change in evaluation practices it will continue to be the case that the first time DNNs are rigorously assessed is after deployment in the real-world. It will therefore continue to be the case that real-world performance will be far lower than expected, and the field will continue to be rightly criticised for making exaggerated claims. Worse leaving rigorous testing until production can have dire consequences, that the research field should feel a moral duty to at least attempt to avoid. The current results show that for the MNIST data-set recent methods of training DNNs produce classifiers that are less robust than ones trained using data augmented with Gaussian noise: a technique that has been around for at least three decades. Without the advocated change in evaluation practices the field is likely to continue to fail to make advances in improving the security, reliability and trustfulness of machine learning models. Worse, lots of time, money, and CO2 (Strubell et al.,, 2020; Thompson et al.,, 2020) will continue to be expended creating methods that give a false sense of progress and security, but when tested thoroughly do not produce improved robustness.
Acknowledgements
The author gratefully acknowledges use of the King’s Computational Research, Engineering and Technology Environment (CREATE).

### v3

@@
 On the other hand, testing with data containing samples from unknown classes fails to evaluate how well the classifier can predict the labels for known classes.
-This article advocates bench-marking performance using a wide range of different types of data and using a single metric that can be applied to all such data types to produce a consistent evaluation of performance.
-Using such a benchmark it is found that current deep neural networks, including those trained with methods that are believed to produce state-of-the-art robustness, are extremely vulnerable to making mistakes on certain types of data.
-This means that such models will be unreliable in real-world scenarios where they may encounter data from many different domains, and that they are insecure as they can easily be fooled into making the wrong decisions.
+This article advocates benchmarking performance using a wide range of different types of data and using a single metric that can be applied to all such data types to produce a consistent evaluation of performance.
+Using the proposed benchmark it is found that current deep neural networks, including those trained with methods that are believed to produce state-of-the-art robustness, are vulnerable to making mistakes on certain types of data.
+This means that such models will be unreliable in real-world scenarios where they may encounter data from many different domains, and that they are insecure as they can be easily fooled into making the wrong decisions.
 It is hoped that these results will motivate the wider adoption of more comprehensive testing methods that will, in turn, lead to the development of more robust machine learning methods in the future.
@@
 To avoid any possible confusion, OOD detection/rejection will hence-forth be referred to as “Unknown Class Rejection”.
-The data used to assess unknown class rejection will be subdivided into “novel” classes (images containing objects from classes unseen during training) and “unrecognisable” images (those containing no sematically meaningful information).
+The data used to assess unknown class rejection will be subdivided (as in Fig.
+1) into “novel” classes (images containing objects from classes unseen during training) and “unrecognisable” images (those containing no sematically meaningful information).
 The term “distribution” when applied to the input is also poorly defined and, arguably, irrelevant when discussing IID/OOD generalisation.
@@
 Furthermore, the shift in distribution in image-space often has little correlation with the difficulty in correctly classifying an image.
-For example, two images of the same object taken from different viewpoints may have very little similarity in terms of corresponding pixels values, but the images may be easily identified as containing the same object.
+For example, two images of the same object taken from different viewpoints may have very little similarity in terms of corresponding pixel values, but the images may be easily identified as containing the same object.
 In contrast, it may be very difficult for the classifier to assign the same label to two images that are almost identical in image-space (e.g., a clean image and an adversarially perturbed version of that same image).
 Most importantly, how far an image is from the training distribution is irrelevant to the task of the classifier.
-Ideally, the classifier should be able to successfully recognise novel exemplars of the objects it was trained on however similar or dissimilar they are in appearance to the images used in training.
+Ideally, the classifier should be able to successfully recognise novel exemplars of the objects it was trained on regardless of how similar or dissimilar they are in appearance to the images used in training.
 In contrast an ideal classifier should reject images containing unknown objects even if such images are close to the training distribution.
-Hence, rather than using the terms IID and OOD, the remaining text will refer to such data as being from “known” classes.
-The broad class of known data will be sub-divided into different types of test data-set: “clean” data from the standard test set, “corrupt” for samples from the standard test set that have been manipulated to test for generalisation to changes in viewing conditions, and “adversarial” for samples that have been adversarially-perturbed.
+Hence, rather than referring to samples as IID and OOD, the remaining text will refer to such data as being from “known” classes.
+The broad class of known data will be sub-divided (as in Fig.
+1) into different types of test data-set: “clean” data from the standard test set, “corrupt” for samples from the standard test set that have been manipulated to test for generalisation to changes in viewing conditions, and “adversarial” for samples that have been adversarially-perturbed.
 The terms “generalisation” and “robustness” both refer to a classifier’s ability to produce the correct predictions.
@@
 In some publications the clean accuracy of the model is evaluated, but this is done without any samples being rejected, and hence, does not reflect the performance that would be expected in the real-world where the same rejection criteria would be applied to all samples.
+As a result of this issue, a model that could accurately reject unknown classes would score highly with the standard metrics, even if that classifier assigned a random label to all accepted samples making it virtually useless in any practical application.
 | predicted class | |||||
@@
 When the threshold is greater or equal to the attack perturbation budget, robustness with rejection is equivalent to standard robust error.
-When the threshold is zero this metric is equivalent to robust risk with detection as proposed by Tramèr, (2021).
+When the threshold is zero this metric is equivalent to robust risk with detection as proposed by Tramèr, (2021), discussed earlier.
 An issue with all the methods discussed in the previous paragraph, is that they define which samples can be rejected in terms of an arbitrary measure of distribution shift in the input space (Guérin et al.,, 2023).
@@
 These different values for the rejection threshold define the trade-off between accuracy on the clean data (perfect performance on the clean data could only be achieved if 100% of correctly classified clean examples were accepted) and correct rejection of unknown samples (good performance on unknown class rejection is more likely if more clean samples are also rejected).
-2 Methods
-2.1 Proposed Assessment Method
-This article proposes to extend and modify the evaluation method proposed by Zhu et al., (2024), described in the preceding section, in the following ways:
+2 Methods: The Proposed Comprehensive Assessment Benchmark
+2.1 Assessment Method
+The proposed benchmark uses a metric to evaluate performance that is an extended and modified version of the evaluation method proposed by Zhu et al., (2024), described in the preceding section.
+The proposed modifications are as follows:
 1.
@@
 Extending the evaluation to consider other forms of generalisation.
-Zhu et al., (2024) evaluate performance using several data-sets including clean and corrupt data and data from unknown classes.
+Zhu et al., (2024) evaluate performance using three types of test data: clean, corrupt, and novel.
 In the current work, unknown class rejection is evaluated not only using images from novel classes, as in (Zhu et al.,, 2024), but also using unrecognisable images.
@@
 1, whereas Zhu et al., (2024) use only three.
+The results reported in Section 3 highlight the necessity of expanding the range of test data types used.
+The poorest performance of the tested networks is almost always for either adversarial samples or unrecognisable images, i.e., for the data-types that Zhu et al., (2024) do not use.
+Hence, the benchmark proposed by Zhu et al., (2024) is more likely to fail to identify robustness issues than the benchmark proposed here.
 2.2 Choice of Test Data
-The current work tested the performance of Deep Neural Networks (DNNs) trained on the following data-sets: CIFAR10, CIFAR100, TinyImageNet, and MNIST.
-In each case training was performed with the standard training sub-set of data.
-The robustness of each trained network was tested using the five categories of test data illustrated in Fig.
+| Training | Test Data | ||||||||||
+| Data-set | Clean | Corrupt | Adversarial | Novel | Unrecognisable | ||||||
+| MNIST | MNIST test-set | MNIST-C | AA | AA | Omniglot | FMNIST | KMNIST | Phase | Scramble | Blobs | Uniform |
+| FMNIST | FMNIST test-set | FMNIST-C | AA | AA | Textures | MNIST | CIFAR10 | ||||
+| SVHN | SVHN test-set | MNIST–C | AA | AA | Textures | Icons50 | CIFAR10 | ||||
+| CIFAR10 | CIFAR10 test-set | CIFAR10-C | AA | AA | Textures | SVHN | CIFAR100 | ||||
+| CIFAR100 | CIFAR100 test-set | CIFAR100-C | AA | AA | Textures | SVHN | CIFAR10 | ||||
+| TinyIN | TinyIN val.-set | TinyIN-C | AA | AA | Textures | iNaturalist | ImageNetO | ||||
+| ImageNet1k | ImageNet1k val.-set | ImageNet-C | AA | AA | Textures | iNaturalist | ImageNetO | ||||
+The benchmark is defined to assess the performance of classifiers trained on seven popular image classification data-sets: MNIST (LeCun et al.,, 1998), FashionMNIST (FMNIST; Xiao et al.,, 2017), SVHN (Netzer et al.,, 2011), CIFAR10 (Krizhevsky,, 2009), CIFAR100 (Krizhevsky,, 2009), TinyImageNet (TinyIN), and ImageNet1k (a.
+k.
+a.
+ILSVRC2012; Russakovsky et al.,, 2015).
+Performance is evaluated using test-data from each of the five categories illustrated in Fig.
 1.
-Within each category the following data was used.
+The specific data-sets used in each category is summarised in Table 2 and described in more detail below.
 - Clean.
@@
 - Corrupt.
-Popular common corruptions data-sets were used: CIFAR10-C, CIFAR100-C, TinyImageNet-C and MNIST-C (Hendrycks and Dietterich,, 2019; Mu and Gilmer,, 2019).
-The first three data-sets contain 18 different corruptions including different types of noise, blurring, synthetic weather conditions, and digital corruptions.
-MNIST-C contains 15 different corruptions including different types of noise, blurring, geometric transformations, and superimposed patterns.
-As typical in the literature, performance was evaluated using all the corruptions at all degrees of intensity.
+Common corruptions data-sets: MNIST-C, FashionMNIST-C (FMNIST-C), CIFAR10-C, CIFAR100-C, TinyImageNet-C and ImageNet-C (Hendrycks and Dietterich,, 2019; Mu and Gilmer,, 2019; Weiss and Tonella,, 2022).
+These data-sets modify the standard test-set using multiple corruption methods including different types of noise, blurring, geometric transformations, superimposed patterns, synthetic weather conditions, and digital corruptions.
+As typical in the literature, performance is evaluated using all the corruptions at all degrees of intensity.
+When the test-set resolution differs from that of the training set, the test images is resized (using bi-linear interpolation) to match the image resolution used for training.
+Similarly, the number of colour channels in the test images is modified, if necessary, to match the training data (e.g., grayscale images are converted to colour by making the three colour channels all equal to the original image intensity values).
 - Adversarial.
-Samples were produced using AutoAttack (AA; Croce and Hein,, 2020), a state-of-the-art ensemble attack method that employs both gradient-based (white-box) and gradient-free (black-box) attacks.
-AA was implemented using the torchattacks PyTorch library (Kim,, 2021).
-Two sets of adversarial samples where created.
-Each set was created by perturbing the entire standard (clean) test-set, but with a different method of constraining the magnitude of the perturbation.
-Specifically, AA was used to apply both and -norm constrained attacks.
-The perturbation budget () used for each attack was the standard value used in the previous literature for each data-set.
-Specifically, was set to and 0.5 for and -norm attacks, respectively, against networks trained on the CIFAR10, CIFAR100, and TinyImageNet data-sets, and was set to 0.3 for -norm and to 2 for -norm attacks on MNIST trained networks.
+Samples are produced using AutoAttack (AA; Croce and Hein,, 2020), a state-of-the-art ensemble attack method that employs both gradient-based (white-box) and gradient-free (black-box) attacks.
+AA is implemented using the torchattacks PyTorch library (Kim,, 2021).
+Two sets of adversarial samples are used.
+Each set is created by perturbing the entire standard (clean) test-setbbbDo to the extremely high computational demands of performing AA on large ImageNet1k trained networks, only 1000 adversarial samples were used for these networks., but with a different method of constraining the magnitude of the perturbation.
+Specifically, AA is used to apply both and -norm constrained attacks.
+The perturbation budget () used for each attack is the standard value used in the previous literature for each data-set, and is defined in Table 2.
 - Novel Class.
-In each case, three data-sets containing objects of unknown classes were chosen.
-Specifically, the CIFAR10 and CIFAR100 trained networks were tested using novel classes from the test-sets of the Textures (Cimpoi et al.,, 2014) and SVHN (Netzer et al.,, 2011) data-sets and the CIFAR data-set that was not used for training (i.e., a CIFAR10 trained network was evaluated on its ability to reject samples from CIFAR100, and vice versa).
-For TinyImageNet trained networks novel class data came from the Textures test set (Cimpoi et al.,, 2014), the iNaturalist 2021 validation set (Van Horn et al.,, 2018), and the ImageNet-O data-set (Hendrycks et al., 2021c, ).
-The MNIST trained networks were tested using unknown, novel, classes from the test-sets of the Omniglot (Lake et al.,, 2015), FashionMNIST (Xiao et al.,, 2017) and KMNIST (Clanuwat et al.,, 2018) data-sets.
+Three data-sets containing objects of unknown classes are used for each training data-set.
+Specifically, MNIST trained networks are tested using unknown, novel, classes from the test-sets of the Omniglot (Lake et al.,, 2015), FashionMNIST (Xiao et al.,, 2017) and KMNIST (Clanuwat et al.,, 2018) data-sets.
+FashionMNIST trained networks are evaluated with the test-sets of the Textures (Cimpoi et al.,, 2014), MNIST and CIFAR10 data-sets.
+When trained with SVHN, unknown class rejection is evaluated with the Textures, Icons-50 (Hendrycks and Dietterich,, 2019), and CIFAR10 test data-sets.
+The CIFAR10 and CIFAR100 trained networks are tested using novel classes from the test-sets of the Textures and SVHN data-sets and the CIFAR data-set that was not used for training (i.e., a CIFAR10 trained network is evaluated on its ability to reject samples from CIFAR100, and vice versa).
+For TinyImageNet and ImageNet1k trained networks novel class data comes from the Textures test set, the iNaturalist 2021 validation set (Van Horn et al.,, 2018), and the ImageNet-O data-set (Hendrycks et al., 2021c, ).
+When the test-set resolution differs from that of the training set, the test images is resized (using bi-linear interpolation) to match the image resolution used for training.
+Similarly, the number of colour channels in the test images is modified, if necessary, to match the training data (e.g., colour images are converted to grayscale).
 - Unrecognisable.
-Data to test unrecognisable image rejection was the same for all networks, and consisted of synthetic images randomly generated using four different methods.
-First, images containing random blobs, as used in (Hendrycks et al., 2019a, ).
-Secondly, images in which each pixel intensity value was independently and randomly selected from a uniform distribution.
-Thirdly, the images of the standard (clean) test set after a random permutation of all pixels.
-Fourthly, the images of the clean test set after randomising the phase, in the Fourier domain, of each image.
-Each of these four data-sets contained the same number of samples as the standard test-set associated with the training data.
+Data to test unrecognisable image rejection is the same for all training data-sets, and consisted of synthetic images randomly generated using four different methods.
+First, the images of the standard (clean) test set after randomising the phase, in the Fourier domain, of each image (“Phase”).
+Second, the images of the clean test set after a random permutation of all pixels (“Scramble”).
+Third, as used in Hendrycks et al., 2019a , images containing random blobs (“Blobs”).
+Fourth, images in which each pixel intensity value is independently and randomly selected from a uniform distribution (“Uniform”).
+Each of these four data-sets contain the same number of samples as the standard test-set associated with the training data.
 In principal, the amount of test data that could be used is unlimited (as there are countless images containing objects that any one classifier has not been trained to recognise, and due to the stochastic nature of the methods used to create corrupt, adversarial, and unrecognisable images).
@@
 In making this selection the aim was to define a reasonably wide variety of data while ensuring that evaluation could be performed in a reasonable time using reasonable resources.
-The results in Section 3 suggest that the chosen data is sufficiently challenging that modern DNNs are far from ceiling performance.
+The results in Section 3 suggest that the chosen data is sufficiently challenging that modern Deep Neural Networks (DNNs) are far from ceiling performance.
 However, as classification methods improve in the future, the data used for evaluation could be extended to allow a more thorough and challenging assessment in two ways: adding more data in each category, or by adding categories.
-Expanding the range of data in each of the existing categories could be done by: using new sets of clean data (such as CIFAR10.1 (Recht et al.,, 2018), or ImageNet-A (Hendrycks et al., 2021b, )); using different corruptions (e.g., different forms of noise or image transformation), higher magnitude corruptions (e.g., a wider range of scales such as used in the MNIST Large Scale data set (Jansson and Lindeberg,, 2020)), the different types of corruption used in the OOD-CV-v2 benchmark (Zhao et al.,, 2024), or objects shown against different backgrounds (as is used in the NICO challenge (Zhang et al., 2023b, )); using any of numerous alternative algorithms for generating adversarial images (Liang et al.,, 2022; Ren et al.,, 2020; Xu et al.,, 2020), or different methods of constraining the perturbations, or higher perturbation budgets; using the additional novel class data-sets proposed in the OpenOOD benchmark Yang et al., (2022); using fooling images (Kumano et al.,, 2022; Nguyen et al.,, 2015) that are unrecognisable images which have been optimised to be difficult to reject.
-Adversarial attack-type techniques could also be used not only to perturb clean test samples, as is currently the case, but samples from other data types to make these types of data more challenging to correctly process.
-New categories of test data that could be used to extend and increase the challenge of the proposed benchmark include: known classes presented in the form of drawings, art, or cartoons (such as in ImageNet-R (Hendrycks and Dietterich,, 2019) or PACS (Li et al.,, 2017)); or images where texture has been replaced by that from another image (Geirhos et al.,, 2019).
+Expanding the range of data in each of the existing categories could be done by: using new sets of clean data, such as CIFAR10.1 (Recht et al.,, 2018), or ImageNet-A (Hendrycks et al., 2021b, ); using different corruptions (e.g., different forms of noise or image transformation), higher magnitude corruptions (e.g., a wider range of scales such as used in the MNIST Large Scale data set (Jansson and Lindeberg,, 2020)), the different types of corruption used in the OOD-CV-v2 benchmark (Zhao et al.,, 2024), or objects shown against different backgrounds (as is used in the NICO challenge (Zhang et al., 2023b, )); using any of numerous alternative algorithms for generating adversarial images (Liang et al.,, 2022; Ren et al.,, 2020; Xu et al.,, 2020), or different methods of constraining the perturbations, or higher perturbation budgets; using the additional novel class data-sets proposed in the OpenOOD benchmark (Yang et al.,, 2022); using fooling images (Kumano et al.,, 2022; Nguyen et al.,, 2015) that are unrecognisable images which have been optimised to be difficult to reject.
+Adversarial attack-type techniques could also be used not only to perturb clean test samples, as is currently the case, but samples from other data types to make these types of data more challenging to correctly process (Li et al.,, 2024).
+New categories of test data that could be used to extend and increase the challenge of the proposed benchmark include: known classes presented in the form of drawings, art, or cartoons, such as in ImageNet-R (Hendrycks and Dietterich,, 2019) or PACS (Li et al.,, 2017); or images where texture has been replaced by that from another image (Geirhos et al.,, 2019).
 2.3 Rejection Criteria
@@
 Many different methods for unknown class rejection have been proposed (Vaze et al.,, 2022; Pang et al.,, 2022; Chen et al.,, 2022; Hendrycks and Gimpel,, 2017; Tajwar et al.,, 2021).
-This article will primarily use maximum softmax probability (MSP; Hendrycks and Gimpel,, 2017), which defines the confidence that a sample is of a known class as the maximum response of the network output after application of the softmax activation function.
+The proposed benchmark uses maximum softmax probability (MSP; Hendrycks and Gimpel,, 2017), which defines the confidence that a sample is of a known class as the maximum response of the network output after application of the softmax activation function.
 MSP is used as this is a standard, baseline method, used frequently in the literature on unknown class rejection.
 Furthermore, a comparison of different rejection methods using the same metric on which the proposed benchmark is based found MSP to be one of the best methods (Zhu et al.,, 2024).
-Previous work has shown that relative performance of different unknown class rejection methods vary depending on a number of factors (Szyc et al.,, 2023) including the data-sets Tajwar et al., (2021) and models (Zhu et al.,, 2024) used.
-Hence, to confirm that the results obtained in this study are not unique to MSP results for three other post-hoc methods of unknown class rejection (those that rely on the responses produced by the network) are also reportedbbbOnly post-hoc methods are considered, as (1) they can be applied to the same network, allowing for a fair comparison and avoiding the need to re-train the network to assess different methods, (2) because post-hoc methods have performance comparable to, or better than, other methods (Yang et al.,, 2022).
-(1) Maximum logit score (MLS; Vaze et al.,, 2022), which defines the confidence that a sample is of a known class as the maximum response of the network output before any activation function is applied.
+Previous work has shown that relative performance of different unknown class rejection methods vary depending on a number of factors (Szyc et al.,, 2023) including the data-sets (Tajwar et al.,, 2021) and models (Zhu et al.,, 2024) used.
+Hence, to confirm that MSP is an appropriate choice, results for three other post-hoc methods of unknown class rejection (those that rely on the responses produced by the network) are also reported in Section 3.1.1: (1) Maximum logit score (MLS; Vaze et al.,, 2022), which defines the confidence that a sample is of a known class as the maximum response of the network output before any activation function is applied.
 This was shown to be the best method of unknown class rejection in a recent comparison of state-of-the-art methods (Vojir et al.,, 2023).
 (2) Energy score (Liu et al.,, 2020), which defines confidence as the negative logarithm of the denominator of the softmax activation function applied the network output layer.
-(3) GEN (Liu et al.,, 2023) a very recently proposed method that claims to be superior to all three other methods considered here.
+(3) GEN (Liu et al.,, 2023) a recently proposed method that claims to be superior to all three other methods considered here.
+Note, only post-hoc methods are considered, as (A) they can be applied to the same network, allowing for a fair comparison and avoiding the need to re-train the network to assess different methods, (B) because post-hoc methods have performance comparable to, or better than, other methods (Yang et al.,, 2022).
 Consistent with the method proposed by Zhu et al., (2024), a rejection threshold is set such that a fixed percentage correctly classified clean examples are accepted.
-In other words, such that the confidence associated with those samples (calculated using either MSP, MLS, Energy score or GEN depending on the experiment being performed), is greater than or equal to the threshold.
+In other words, such that the confidence associated with those samples, is greater than or equal to the threshold.
 All samples, from any data-set, with confidence lower than the threshold are rejected, while all others are accepted and the class label predicted.
-2.4 Choice of Models to Assess
-The aim of the current work is to determine how well existing DNNs perform when evaluated using the more comprehensive assessment methods advocated above.
-However, it is clearly impossible to attempt to evaluate the performance of every network architecture or to evaluate the huge variety of different techniques that has been proposed for improving adversarial robustness, generalisation to common corruptions, or for unknown class rejection.
-ResNet18 (He et al.,, 2016) was chosen as the principal DNN architecture to test as this is very commonly used in the literature.
-This model has just over 11M trainable parameters.
-Additional experiments were also performed using WideResNet34-20 (Zagoruyko and Komodakis,, 2016) which has just under 193M parameters.
-For the MNIST data-set smaller networks were used.
-Firstly, a full-connected multi-layer perceptron network, containing 200 neurons in each of 3 hidden layers (denoted as MLP200-200-200).
-This network used the ReLU activation function and had 239,410 trainable parameters.
-Secondly, a fully-convolutional neural network (ConvNet) containing 5 layers, with 32 3x3 convolutional filters in each hidden layer, and average pooling of stride 2 applied after the first and second layers.
-The ReLU activation function was applied to each layer (except the final output).
-This network is denoted as ConvNet5-3-32 where the numbers indicate the depth, kernel-size, and width respectively.
-This network contained 30,954 trainable parameters.
-A common approach to improving performance on all the different types of test-data described in Section 1.2 is to use data augmentation techniques: i.e., to expand the training data-set by creating new samples that are synthetically manipulated versions of samples from the original training data-set.
-In the results section (Section 3) a number of such techniques are evaluated.
-The selection of the particular methods to test was made so as to include standard baselines and state-of-the-art methods for producing robust models for all the different types of test-data described in Section 1.2.
-- baseline
-Improved performance on the standard test data is often achieved by applying simple geometric data augmentations to the training data.
-Here 4-pixel random cropping and random horizontal flipping where applied to the CIFAR10, CIFAR100 and TinyImageNet training data, while no augmentations were applied to the MNIST data.
-These standard augmentations were used in combination with the additional augmentation methods listed below.
-Given that these baseline augmentation methods are very commonly used when training models, the results show how standardly trained networks cope under comprehensive evaluation.
-The results for this method of training also provides a baseline against which to judge the effectiveness of other methods that have been proposed to improve generalisation performance.
-- noise
-Augmenting training images with random noise has been shown to reduce over-fitting, and hence, improve generalisation (Sietsma and Dow,, 1991; Holmstrom and Koistinen,, 1992; Goodfellow et al.,, 2016), especially to common image corruptions (Rusak et al.,, 2020; Gilmer et al.,, 2019).
-Various methods of adding noise to the training data have been proposed.
-Rusak et al., (2020) augmented half the training data with Gaussian noise, and found that this technique produced performance on corrupt test data that was competitive with the more complex augmentation method, AugMix, and that for the MNIST data-set specifically was just as effective as their proposed method of Adversarial Noise Training.
-Lopes et al., (2019) advocated adding Gaussian noise to only a randomly chosen patch of the image, which they claimed produced better robustness to image corruptions on CIFAR10 data than applying random noise to the whole image.
-Gilmer et al., (2019) used an alternative method for corrupting training images with Gaussian noise in which the standard deviation of the noise applied to each image was chosen with equal probability from the range .
-The method implemented here is a generalisation for the method proposed by Gilmer et al., (2019).
-Specifically, for each image the standard deviation of the noise was chosen from a uniform distribution with range .
-Then each pixel intensity value was independently modified by the addition of a value chosen at random from a Gaussian distribution with mean of zero and the chosen standard deviation.
-The original images had pixel intensity values in the range [0,1], and after the application of the noise pixel values were clipped at 0 and 1 to remain in this range.
-Appropriate values of and were found via a grid search using small ConvNets.
-Using this method and were chosen for use with CIFAR10, and the same hyper-parameters were also used for CIFAR100 and TinyImageNet.
-For MNIST was used.
-- AT
-Multi-step adversarial training with Projected Gradient Descent (PGD; Madry et al.,, 2018) is a standard, and highly effective, defence against adversarial attack.
-As a result, this method has become a standard benchmark against which all other methods of adversarial defence are judged.
-This method was, therefore, also selected as a data augmentation technique to be tested.
-It was implemented using the torchattacks PyTorch library (Kim,, 2021).
-To perform adversarial training, clean images from the current batch of training examples were used to generate non-targeted adversarial examples, that then replaced the original training images (as in Madry et al.,, 2018).
-Following standard protocols, the maximum allowed perturbation was constrained by the -norm to be less than the perturbation budget (), and the training-time adversarial images were generated using 10 steps of PGD (denoted as PGD for brevity).
-Standard values used in the previous literature were used for .
-Specifically, was set to for performing AT with the CIFAR10, CIFAR100 and TinyImageNet training data, and was used for MNIST.
-- pixmix
-This is a training data augmentation pipeline that repeatedly mixes (either additively or multiplicatively) an augmented image with an augmented clean image or a fractal image (Hendrycks et al.,, 2022).
-It is claimed that this augmentation method can improve generalisation to both clean and corrupt data, improve adversarial robustness, and improve unknown class rejection.
-This method of data augmentation was therefore also tested using the proposed, comprehensive, evaluation method.
-It was implemented using the official code released by the original authors.
-- regmixup
-In addition, evaluations were also performed on current state-of-the-art robust pre-trained models available from the RobustBench model zooccchttps://github.com/RobustBench/robustbench (Croce et al.,, 2021).
-2.5 Experimental Setup
-Training was performed using a batch size of 128.
-For the CIFAR10, CIFAR100 and TinyImageNet data-sets, training was performed for 110 epochs using the SGD optimiser with weight decay of 5e-4 and an initial learning rate of 0.1.
-The learning rate was step-wise decayed by a factor of 10 at the start of epochs 100 and 105 for the CIFAR data-sets, and at epochs 45 and 89 for TinyImageNet.
-The set-up for the CIFAR data-sets was based on that recommended for AT by Pang et al., (2021).
-Hence, these hyper-parameters may not be optimal for training with the other data augmentations used.
-For the MNIST data, 20 epochs of training using the Adam optimiser and a constant learning rate of 0.001 was used.
-For each experimental condition (combination of data-set, network architecture, and training data augmentation method) the network was trained and evaluated three times, each time with a different random weight initialisation and random presentation order of training samples.
-The reported result for each condition is the mean accuracy over these three trials.
-The standard-deviation in the mean performance measured using the proposed metric is shown in the tables of results and generally shows a high degree of consistency across trials.
-Exceptions to the above, where only a single trial was performed, were the experiments with WideResNet34-20 and the experiments performed using a publicly available checkpoint created by other researchers.
-Open-source code written in Python using the Pytorch library (Paszke et al.,, 2019), which performs all the experiments described in this article, is available for download from: https://codeberg.org/mwspratling/RobustnessEvaluation.
+Zhu et al., (2024) propose setting the threshold such that either 95% or 99% of correctly classified clean examples are accepted.
+For the proposed benchmark 95% is used, but results for the alternative are also reported in Section 3.1.2.
+Instead of using a fixed threshold, it would also be possible to calculate performance over a range of thresholds, as with the AUROC and AUPR metrics.
+However, in a practical application it would be necessary to define a fixed threshold, so the proposed approach is more likely to accurately reflect real-world performance.
+Furthermore, the creator of the classifier is likely to want it to perform well on the data it was trained on, so will want to accept a high proportion of the correctly classified clean samples.
 3 Results
-3.1 Comparison with Standard Metrics
+Section 3.1 reports results of experiments designed to validate the proposed benchmark and justify some of the choices that have been made in defining it (Sections 3.1.1 and 3.1.2).
+Section 3.2 reports results designed to determine how well existing DNNs perform when evaluated using the more comprehensive assessment methods being advocated.
+The proposed benchmark can be used to assess the performance of any classifier trained on any of the supported data-sets.
+Obviously, it would be impossible in this one publication to attempt to benchmark the performance of every such classifier or to evaluate the huge variety of different techniques that can be used with such classifiers to improve adversarial robustness, generalisation performance, or unknown classes rejection.
+Hence, results are only reported for a few, select, models.
+The first set of experimental results (Section 3.1) were produced using self-trained ResNet18 models (He et al.,, 2016).
+These models were trained on CIFAR10 and CIFAR100 using a variety of training-time data augmentations that have previously been proposed to improve robustness or generalisation.
+Augmentation methods that produce different levels of robustness to different test data types were chosen, so as to be able to ensure that this variation in performance is captured by the proposed benchmark.
+These networks have been trained using standard procedures full details of which are given in Appendix A.
+The second set of results (Section 3.2) focus of benchmarking the performance of state-of-the-art (SOTA) robust models.
+These models were taken from the RobustBench leaderboards.
+RobustBench (Croce et al.,, 2021) is an existing benchmark for assessing the robustness of DNNs.
+Unlike the proposed benchmark, RobustBench evaluates performance of a model using only a single type of data: either adversarial (produced using either or constrained AA) or common corruptions.
+From each of these RobustBench leaderboards a selection of the top performing models were evaluated using the proposed benchmark.
+The selected models included the highest ranked model from each leaderboard for which pre-trained weights were available from the RobustBench repositoryccchttps://github.com/RobustBench/robustbench on November 15th 2024.
+Other tested models were selected from the top-10 with a bias towards the highest ranked models, but lower ranked models were sometimes selected to increase the diversity of network architectures and robustification methods evaluated.
+3.1 Experiments to Validate the Benchmark
+This paper advocates for (1) more comprehensive evaluation of DNN performance (Section 1.4), and (2) using a single, consistent, metric to assess performance on all types of testing data (Section 1.5).
+Figs.
+3(a), 3(a), 3(a) and 3(e) shows performance for many different test data sets, measured using existing metrics.
+Hence, these results address aim (1) but not aim (2).
+As discussed in Section 1.4 such comprehensive assessment is rarely performed.
+However, it provides valuable insights into both the strengths and weaknesses of different classifiers.
+For example, it can be seen that the networks trained using adversarial training (AT) with PGD are most resistant to adversarial attacks with AA, but that such networks have reduced performance on unknown class rejection (and are particularly vulnerable to making incorrect classifications when presented with images containing scrambled pixels).
+The former is expected, but the latter is ignored in the adversarial robustness literature as evaluations with such data are never performed.
+Similarly, it is to be expected given previous research that pixmix produces excellent performance on unknown class rejection, however, the vulnerability of such networks to adversarial attack has been mostly undocumented.
 | Training | Clean | Corrupt | AA | AA | Textures | SVHN | CIFAR100 | Phase | Scramble | Blobs | Uniform |
-| Method | Accuracy (%) | AUROC (%) | |||||||||
+| Method | Accuracy (%) | AUROC (%) when MSP used as prediction confidence score | |||||||||
 | baseline | |||||||||||
 | noise | |||||||||||
-[…]
-[…]
-[…]
-| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
-| baseline | .77) | |||||
-| noise | .99) | |||||
 […]
@@
 […]
-| Training Method | Clean | Corrupt | AA | AA | Textures | SVHN | CIFAR10 | Phase | Scramble | Blobs | Uniform |
-| Accuracy (%) | AUROC (%) | ||||||||||
-[…]
-| noise | |||||||||||
-[…]
-[…]
-| regmixup | |||||||||||
-| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
-[…]
-| noise | .45) | |||||
-[…]
-[…]
-| regmixup | .28) |
-| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
-[…]
-| noise | .09) | |||||
-[…]
-[…]
-| regmixup | .97) |
-Table 2 shows results for ResNet18 networks trained on CIFAR10 using different training-time data augmentation methods that have been proposed to improve performance (as discuss in Section 2.4).
-Table 3 shows equivalent results for CIFAR100 trained networks.
-Evaluation of the performance of these networks using standard metrics is shown in Figs.
-3(a) and 3(d).
-It can be seen that, as expected, AT with PGD produces the best defence against adversarial attacks with AA.
-It can also be seen that, compared to a network trained with the baseline augmentations, AT causes a reduction in the accuracy with which clean data is categorised and a reduction in performance on unknown class rejection.
-Also not unexpectedly, pixmix produces the best performance on unknown class rejection for all the data-sets (except CIFAR10 for the CIFAR100 trained network), and produces the best performance on the corrupt data.
-These findings are also captured by the new evaluation metric, results for which are shown in Figs.
-3(b) and 3(e) when the rejection threshold is set such that 99% of correctly classified clean examples are accepted.
-The results summarising performance on each type of data show that AT produces the highest accuracy for adversarial samples, and pixmix produces the highest accuracy on corrupt data and the two types of unknown class data (novel and unrecognisable).
-The trade-offs in performance over different types of data can also be seen.
-The average performance over all data types (shown in the last columns of each table) suggests that pixmix produces the best performance overall.
-The results for all methods on unknown class data is far worse than the standard metrics would imply.
-This is because the same rejection threshold has been used for all data types, rather than being tuned to optimise performance on unknown class rejection.
-This illustrates that current metrics for unknown class rejection grossly over-estimate likely real-world performance (as discussed in Section 1.5).
-In contrast, the performance of all methods on known class data (clean, corrupt and adversarial) is higher than would be expected from the standard metrics.
-This is because networks can reject samples for which prediction confidence is low which include many samples for which the predicted class is wrong.
-Note that these results generally show that current methods of training deep networks produce classifiers that lack robustness.
-For most of the methods evaluated (which includes methods that are designed to produce state-of-the-art robustness) the resulting network can only make the correct decision on around 50% of the samples for CIFAR10, which is a relatively simple data-set.
-Even pixmix, that performs best overall, only makes accurate decisions for just over 71% of the test data.
-Furthermore, it might reasonably be argued that a model is only as robust as the worst results would suggest, as a malicious actor can select the type of data used to attack a system, and in the real world a model might naturally encounter all types of data and would need to deal correctly with all.
-For each method of training investigated here there is at least one type of test data on which the model performs pitifully.
-For example, AT can only correctly reject 2.2% of unrecognisable images: all the others are given incorrect class labels with a high confidence.
-Massive amounts of time and resources have gone into developing techniques to enhance adversarial robustness.
-The current results suggest that this effort may have been wasted, as an attacker can easily and effectively induce a classifier to generate the wrong predictions by presenting it with easily generated synthetic images.
-The results in Section 3.5 show that the most recent state-of-the-art methods of adversarial robustness, while having better performance on unknown class rejection, are still far from robust to such data.
-Figs.
-3(c) and 3(f) show results for the proposed evaluation method when the rejection threshold is set such that 95% of correctly classified clean examples are accepted (an increased rejection threshold compared to previously).
-For the CIFAR10 trained networks, rejecting more of the clean samples results in a reduction in performance on the clean data as some samples that were correctly classified have been rejected.
-In contrast, the results for CIFAR100 show that performance on the clean data improves with the increased rejection threshold, due to the higher proportion of clean samples that are mis-classified by the CIFAR100 trained networks.
-For both data-sets, rejecting more samples for which confidence in the prediction is low results in a large improvement in performance on the adversarial, novel and unrecognisable data types for all training methods.
-The results produced with both thresholds are consistent showing similar trade-offs in performance on the different data types for each of the tested methods, and producing similar rankings of different training methods, although the differences between methods is somewhat reduced when the threshold is set so as to accept 95% of correctly classified clean samples.
-To avoid redundancy, and for brevity, only results for one threshold are reported subsequently.
-The 95% threshold was chosen in order to present the most favourable performance for the tested DNNs.
-3.2 Different Unknown Class Rejection Methods
 | Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std | |
@@
 | regmixup | .15) |
+| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
+| baseline | .77) | |||||
+| noise | .99) | |||||
+[…]
+[…]
+[…]
+| Training Method | Clean | Corrupt | AA | AA | Textures | SVHN | CIFAR10 | Phase | Scramble | Blobs | Uniform |
+| Accuracy (%) | AUROC (%) when MSP used as prediction confidence score | ||||||||||
+[…]
+| noise | |||||||||||
+[…]
+[…]
+| regmixup | |||||||||||
+| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
+[…]
+| noise | .09) | |||||
+[…]
+[…]
+| regmixup | .97) |
 | Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std | |
@@
 | regmixup | .76) |
-The same networks evaluated in the previous section were re-evaluated using MLS, Energy score and GEN as the criteria for rejection (see Section 2.3).
-The results for the networks trained on CIFAR10 are shown in Table 4, and for CIFAR100 in Table 5.
-Comparing the results in Table 4 with the results obtained using MSP shown in Fig.
-3(c) it can be seen that while individual accuracy values can vary, a similar overall performance is produced using each method of rejection, and as a result a similar ranking of the overall performance of each training method is obtained.
-The same can be said for the CIFAR100 results shown in Table 5 and Fig.
-3(f).
-The results obtained, therefore, do not seem to be specific to the particular method of rejection used, and to avoid redundancy, and for brevity, only results for MSP are reported subsequently.
-This choice was made as the worse performance on any data type is better for MSP than for any of the other three methods, and hence, MSP seems to have a slight advantage in improving robustness.
-| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean |
+| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
+[…]
+| noise | .45) | |||||
 […]
 […]
+| regmixup | .28) |
+Figs.
+3(b) and 3(f) show the assessment of the same networks produced using the proposed benchmark.
+Hence, these results address both concerns (1) and (2) described in Sections 1.4 and 1.5.
+The new evaluation metric successfully captures the same pattern of relative performance that can be seen when using the standard metrics.
+For example, AT produces the highest DAR for adversarial samples, but this improvement comes at the cost of reduced performance (compared to the baseline) in clean accuracy and novel and unrecognisable image rejection.
+Of the tested augmentation methods, pixmix produces the highest DAR on corrupt data and the two types of unknown class data, which is consistent with the results produced using the standard metrics.
+regmixup produces the highest performance on clean test data, a result which is captured by both the proposed and standard metrics.
+The proposed metrics differ from the standard ones in some respects.
+For example, for all the tested data augmentation techniques, performance on unknown class data is far worse when measured using DAR than the standard metrics would imply.
+This is because the same rejection threshold has been used for all data types, rather than being tuned to optimise performance on unknown class rejection.
+This illustrates that current metrics for unknown class rejection grossly over-estimate likely real-world performance (as discussed in Section 1.5).
+In contrast, the performance of all the tested models on known class data (clean, corrupt and adversarial) is generally higher than would be expected from the standard metrics.
+This is because networks can reject samples for which prediction confidence is low which include many samples for which the predicted class is wrong.
+An advantage of the proposed benchmark is the use of consistent units to measure performance on each type of test data.
+This facilitates the comparison of performance of different data types.
+Using a consistent performance metric for each test data type also allows an average to be calculated.
+In contrast, it would be meaningless to average the standard metrics given that they are reported in different units.
+The average performance over all data types (shown in the last columns of Figs.
+3(b) and 3(f)) suggests that, of the tested data augmentation methods, pixmix produces the best performance overall.
+While the average performance helps to identify the most promising methods of improving robustness, the individual results for each test data type help identify remaining vulnerabilities that it would most useful to address in future research.
+3.1.1 Choice of Unknown Class Rejection Method
+The proposed benchmark uses MSP as the confidence score used to decided if samples are accepted or rejected.
+The same networks evaluated in the previous section were re-evaluated using alternative rejection methods: MLS, Energy score and GEN (see Section 2.3).
+The results for the networks trained on CIFAR10 are shown in Fig.
+3(c), and for CIFAR100 in Fig.
+3(g).
+Comparing the results in Fig.
+3(c) with the results obtained using MSP shown in Fig.
+3(b) it can be seen that while individual accuracy values can vary, a similar overall performance is produced using each method of rejection, and as a result a similar ranking of the overall performance of each training method is obtained.
+The same can be said for the CIFAR100 results shown in Fig.
+3(g) and Fig.
+3(f).
+The benchmark’s results, therefore, do not seem to be specific to the particular method of rejection used.
+Furthermore, the choice of MSP can be justified as the worse performance on any data type is better for MSP than for any of the other three methods, and hence, MSP seems to have a slight advantage in improving robustness.
+3.1.2 Choice of Unknown Class Rejection Threshold
+The proposed benchmark sets the rejection threshold such that 95% of correctly classified clean examples are accepted (see Section 2.3).
+Figs.
+3(d) and 3(h) shows results for the same networks tested earlier but using a threshold set so that 99% of correctly classified clean examples are accepted.
+The results produced with both thresholds are consistent showing similar trade-offs in performance on the different test data types, and producing similar rankings of different training methods.
+The benchmark’s results, therefore, do not seem to be specific to the particular rejection threshold used.
+In general setting the threshold so that 95% of correctly classified clean examples are accepted produces more favourable evaluation of each classifier.
+This is to be expected, as lowering the percentage of correctly classified clean examples that are accepted, increases the rejection threshold and allows more samples associated with low confidence predictions to be rejected.
+Such low confidence prediction are generally more likely to be wrong, so rejecting them is beneficial.
+3.2 Experiments to Evaluate Model Robustness
+The results in this section aim to determine how well existing DNNs perform when evaluated using the more comprehensive assessment methods advocated in this paper.
+Classifiers with SOTA robustness, as determined by RobustBench (Croce et al.,, 2021), are evaluated and the results are shown in Table 5.
+The comprehensive assessment provides insights into these models that are not unavailable with the narrower analysis provided by RobustBench.
+For example, the CIFAR10 and CIFAR100 models that have been trained to produce good performance on the common corruptions data have surprisingly good performance at unknown class rejection (Figs.
+3(i) and 3(j), upper rows).
+Hence, despite their relative susceptibility to adversarial attack, their overall performance is competitive with CIFAR10/100 models that have been trained for adversarial robustness.
+The CIFAR10 and CIFAR100 models ranked highly by RobustBench for robustness to constrained adversarial attacks are generally very poor at rejecting unrecognisable images (Figs.
+3(i) and 3(j), lower rows).
+As a specific example, RobustBench’s top performing model on CIFAR100 adversarial robustness (Wang et al.,, 2023) produces the wrong predictions around 94% of the time on such images.
+Hence, these models lack the robustness that conventional methods of evaluation suggest they have.
+On the contrary, the proposed comprehensive evaluation shows that SOTA methods of improving performance against adversarial attack fail to provide security as an attacker can use cheaply and easily generated random images to fool the classifier with a high chance of success.
+The CIFAR10 models developed to have robustness to constrained adversarial attacks, compared to their counterparts trained for robustness, are much better at dealing correctly with samples from unknown classes (Fig.
+3(i), middle rows).
+This, perhaps, suggests that diversifying the methods used to produce training-time adversarial images might help reduce this particular vulnerability.
+Another approach to improving performance is also suggested by the performance of the XCiT-L12 model (Debenedetti et al.,, 2023) (Fig.
+3(j), last row).
+This is the only transformer architecture to make it into the top-10 of any of the CIFAR10/100 RobustBench leaderboards (at ninth place for CIFAR100 adversarial robustness).
+This model has far better performance at rejecting unrecognisable images, suggesting that architectural changes may be helpful.
+However, this improvement in robustness to unrecognisable images comes at the cost of increased vulnerability to novel images.
+As a result, an attacker could present this model with random images of objects with classes that differ from those in the training data-set and the model would fail 66% of the time.
+Overall, the proposed comprehensive assessment benchmarks ranks this model as the most robust of the CIFAR100 models tested: a very different ranking than that produced by the narrow testing performed in RobustBench.
+It is remarkable how little improvement in overall, average, performance is shown by the SOTA CIFAR10 and CIFAR100 models (Figs.
+3(i) and 3(j)) compared with the corresponding results in Figs.
+3(b) and 3(f) for (generally) much simpler models made robust with (generally) older methods.
+Indeed, the best performing CIFAR100 RobustBench model has worse performance, on average, than the best performing model in Fig.
+3(f).
+This is due to the SOTA models failing to make progress on improving performance on all the different types of test data.
+For example, if we compare the models trained to be adversarially robust (bottom three rows of Fig.
+3(j) and the row labelled AT in Fig.
+3(f)) it can be seen that the RobustBench models have substantially improved performance on adversarial and clean images, but (as noted above) they remain poor at correctly rejecting samples from unknown classes.
+Massive amounts of time and resources have gone into developing techniques to enhance adversarial robustness beyond that produced by AT with PGD (Madry et al.,, 2018).
+The current results suggest that much of this effort may have been wasted, as these models remain vulnerable as an attacker can easily and effectively induce a classifier to generate the wrong predictions by presenting images from unknown classes.
+The performance of ImageNet1k trained models (Fig.
+3(k)) compared with that of the CIFAR10 models (Fig.
+3(i)) suggests that the benefits of an increased training data-set size is off-set by the increase in task difficulty, so that the performance of the ImageNet1k models is no better than that of the CIFAR10 ones.
+The ImageNet1k results, and all the others that have been presented here, show that current methods of training deep networks produce classifiers that lack robustness.
+In the real world a model might naturally encounter all the diverse types of data that are used in the proposed comprehensive benchmark and would need to deal correctly with all of them.
+However, the best performing models would only make accurate decisions for just over 77% of such data.
+Furthermore, for each of the models evaluated here there is at least one type of test data on which the model performs poorly, which could be exploited by a malicious actor to reduce performance even further.
+| Model | Params. | Robustness | Clean | Corrupt | Adv. | Novel | Unrecog. | Mean |
 […]
+| Diffenderfer et al., (2021) | ||||||||
 […]
-| regmixup |
-| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean |
-| baseline | ||||||
-| noise | ||||||
+| Kireev et al., (2022) | ||||||||
 […]
+| Wang et al., (2023) | ||||||||
 […]
+| Rebuffi et al., (2021) | ||||||||
+| WideResNet94-16 | 366M | adv. () | 92.04 | 86.72 | 86.36 | 59.26 | 21.21 | 69.12 |
+| Bartoldson et al., (2024) | ||||||||
 […]
-| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
-| baseline | .00) | |||||
-| noise | .01) | |||||
+| Amini et al., (2024) |
+| Model | Params. | Robustness | Clean | Corrupt | Adv. | Novel | Unrecog. | Mean |
 […]
+| Diffenderfer et al., (2021) | ||||||||
 […]
+| Modas et al., (2022) | ||||||||
+| ResNeXt29 32x4d | 6.90M | corruptions | 83.71 | 74.80 | 9.04 | 39.68 | 28.56 | 47.16 |
+| Hendrycks et al., (2020) | ||||||||
 […]
-| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
-| baseline | .12) | |||||
+| Wang et al., (2023) | ||||||||
+| WideResNet28-10 | 36M | adv. () | 80.12 | 73.62 | 62.32 | 45.58 | 7.96 | 53.92 |
+| Cui et al., (2024) | ||||||||
 […]
-| AT | .21) | |||||
+| Debenedetti et al., (2023) |
+| Model | Params. | Robustness | Clean | Corrupt | Adv. | Novel | Unrecog. | Mean |
 […]
+| Tian et al., (2022) | ||||||||
+| ResNet50 | 25.6M | corruptions | 80.90 | 74.77 | 15.75 | 37.83 | 58.85 | 53.62 |
+| Hendrycks et al., 2021a | ||||||||
 […]
-| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean Std |
-| baseline | .30) | |||||
+| Liu et al., (2025) | ||||||||
 […]
-[…]
-[…]
-[…]
-3.3 Effects of Increasing Network Capacity
-An increase in network capacity is typically accompanied by an increase in performance.
-Hence, in an attempt to improve on the results presented in Section 3.1 these experiments were repeated using the WideResNet34-20 architecture, that has more than 17 times the number of trainable parameters compared with the ResNet18 architecture used previously.
-The results for CIFAR10 are shown in Table 6.
-Comparing these results to those produced under equivalent conditions with ResNet18 shown in Fig.
-3(c), it can be seen that, surprisingly, results for regmixup are worse for the network with higher capacity, due to a poorer ability to reject samples from unknown classes.
-In contrast, the overall performance improves when baseline, noise, AT, and pixmix are used as the training data augmentation methods.
-The improvement for pixmix is only very modest, while that of AT is the largest.
-The large improvement in the overall performace for AT is accompanied by a large increase in robustness to unrecognisable data.
-However, the results on WideResNet34-20 show that for this training method, as for all the others, there is at least one type of test data on which the model performs poorly.
-Comparing the results for CIFAR100 in Table 7 and Fig.
-3(f) it can be seen the the performance produced by WideResNet34-20 is superior to that produced by ResNet18 for all the different training methods, and the largest increase in performance is for regmixup.
-However, as for the CIFAR10 results, even with a large capacity network the robustness of the networks produced by each training method is still poor and there is still at least one type of test data for which performance is very poor.
-3.4 Results for Other Data-sets
-The determine how well the preceding results generalise, experiments were repeated using a more challenging data-set (TinyImageNet) and an easier one (MNIST).
-For MNIST different network architectures were also used.
-The results for TinyImageNet (Table 8) are very similar to the equivalent results produced for CIFAR100 (Fig.
-3(f)), but with a small drop in overall performance for each training method.
-In contrast, the results for MNIST (Tables 9 and 10) differ from the results for the other data-sets and differ in details between the two architectures used.
-Specifically, the best overall robustness is obtained for the MNIST trained networks, as this is the least challenging data-set.
-However, surprisingly the best performance is not produced by pixmix, as for the other data-sets, but by training with noise.
-3.5 Benchmarking Leading Methods from RobustBench
-| Model | Robustness | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean |
-[…]
-[…]
-| Bartoldson et al., (2024) | adv. () | ||||||
-[…]
-[…]
-[…]
-| Model | Robustness | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean |
-[…]
-[…]
-[…]
-| Cui et al., (2024) | adv. () |
-RobustBench (Croce et al.,, 2021) is an existing benchmark for assessing the robustness of DNNs.
-Unlike the proposed benchmark, RobustBench evaluates performance of a model using only a single type of data: either adversarial (produced using either or constrained AA) or common corruptions.
-The two best performing models from each RobustBench category (for which pretrained models were available from the RobustBench repository on November 15th 2024) were evaluated using the proposed benchmarkdddAA crashed when evaluating the second best method for CIFAR100 adversarial robustness (Bai et al.,, 2024) due to the model producing NaNs.
-This was the case when testing using AA implemented using the torchattacks library (as has been used to adversarially attack all other methods assessed in this paper), and when using the autoattack library’s version of AA (the implementation used by RobustBench).
-Results for the first and third best methods are therefore presented for CIFAR100 adversarial robustness..
-The results are shown in Tables 11 and 12 for the CIFAR10 and CIFAR100 data-sets, respectively.
-Note that these methods use the following model architectures: WideResNet18-2 (268M parameters) (Diffenderfer et al.,, 2021); ResNet18 (11M parameters) (Kireev et al.,, 2022; Modas et al.,, 2022); WideResNet94-16 (366M parameters) (Bartoldson et al.,, 2024); MeanSparse RaWideResNet70-16 (267M parameters) (Amini et al.,, 2024); WideResNet70-16 (267M parameters) (Wang et al.,, 2023; Rebuffi et al.,, 2021); WideResNet28-10 (36M parameters) (Cui et al.,, 2024).
-It can be seen that the state-of-the-art RobustBench methods suffer from the same issues observed in the previous experiments.
-For the models trained to generalise to common corruptions the performance against adversarial attacks is poor.
-While the models trained to be robust to adversarial attacks are poor at rejecting unrecognisable images.
-Hence, overall, these models lack the robustness that conventional methods of evaluation suggest they have.
-On the contrary, the proposed comprehensive evaluation shows that methods of improving performance against adversarial attack, for example, fail to provide security as an attacker can use cheaply and easily generated random images to fool the classifier with a high chance of success.
+| Singh et al., (2023) |
 4 Discussion
@@
 This article promotes an existing metric (Zhu et al.,, 2024) that is capable of being used with different types of data to provide a consistent and comprehensive evaluation of performance.
-Using the proposed benchmarkeeeNote that the proposed evaluation benchmark is likely to be imperfect, and is likely to be biased: like all assessment methods it will favour certain models over others, those that are good at processing the specific data-sets that has been chosen for performing the evaluation, but may be poorer at dealing with other data.
+The proposed benchmark is likely to still be imperfect, and is likely to be biased: like all assessment methods it will favour certain models over others, those that are good at processing the specific data-sets that has been chosen for performing the evaluation, but may be poorer at dealing with other data.
 However, the proposed benchmark is less biased than those that make no attempt to evaluate robustness to certain types of data, and (as discussed in Section 2.2) the data sets used for assessment in the proposed benchmark can easily be extended.
-it is found that existing methods of training DNNs, including methods that are claimed to produce state-of-the-art robustness, fail to perform as well as previous, less comprehensive, assessments would suggest.
+Using the proposed benchmark it is found that existing methods of training DNNs, including methods that are claimed to produce state-of-the-art robustness, fail to perform as well as previous, less comprehensive, assessments would suggest.
 Furthermore, it appears that there are trade-offs that mean that training methods that have good performance on one type of data (i.e., improve the placement of the decision boundary in one region of feature-space) suffer from poor performance on other data (i.e., place the decision boundary less accurately elsewhere in feature-space).
-This means that all methods of training DNNs that have been tested here are highly susceptible to malicious attacks and, by appropriate choice of data type, can be easily fooled into producing the wrong predictions.
-Improved performance might be possible by combining current methods: e.g., using one model to perform unknown class rejection, and an adversarially robust model to classifying the unrejected samples.
+This means that all methods of training DNNs that have been tested here are highly susceptible to malicious attacks and, by appropriate choice of data type, can be fooled into producing the wrong predictions with high probability.
+Improved performance might be possible by combining current methods: e.g., using one model to perform unknown class rejection, and an adversarially robust model to classify the unrejected samples.
 However, appropriate evaluation protocols, such as the one advocated here, will still be needed to test such ensemble methods.
@@
 Without such a change in evaluation practices it will continue to be the case that the first time DNNs are rigorously assessed is after deployment in the real-world.
-It will therefore continue to be the case that real-world performance will be far lower than expected, and the field will continue to be rightly criticised for making exaggerated claims.
+Real-world performance will be far lower than expected, and the field will continue to be rightly criticised for making exaggerated claims.
 Worse leaving rigorous testing until production can have dire consequences, that the research field should feel a moral duty to at least attempt to avoid.
-The current results show that for the MNIST data-set recent methods of training DNNs produce classifiers that are less robust than ones trained using data augmented with Gaussian noise: a technique that has been around for at least three decades.
 Without the advocated change in evaluation practices the field is likely to continue to fail to make advances in improving the security, reliability and trustfulness of machine learning models.
-Worse, lots of time, money, and CO2 (Strubell et al.,, 2020; Thompson et al.,, 2020) will continue to be expended creating methods that give a false sense of progress and security, but when tested thoroughly do not produce improved robustness.
+Worse, lots of time, money, and CO2 (Strubell et al.,, 2020; Thompson et al.,, 2020) will continue to be expended creating methods that give a false sense of progress and security, but when tested thoroughly do not produce substantially improved robustness.
 Acknowledgements
 The author gratefully acknowledges use of the King’s Computational Research, Engineering and Technology Environment (CREATE).
+Appendix A Appendix: Methods used to Train Models
+A.1 Training-Time Data Augmentations
+A common approach to improving performance on all the different types of test-data described in Section 1.2 is to use data augmentation techniques: i.e., to expand the training data-set by creating new samples that are synthetically manipulated versions of samples from the original training data-set.
+In the results section (Section 3) a number of such techniques are evaluated.
+The selection of the particular methods to test was made so as to include standard baselines and state-of-the-art methods for producing robust models for all the different types of test-data described in Section 1.2.
+- baseline
+Improved performance on the standard test data is often achieved by applying simple geometric data augmentations to the training data.
+Here 4-pixel random cropping and random horizontal flipping where applied to the CIFAR10, CIFAR100 and TinyIN training data, while no augmentations were applied to the MNIST data.
+These standard augmentations were used in combination with the additional augmentation methods listed below.
+Given that these baseline augmentation methods are very commonly used when training models, the results show how standardly trained networks cope under comprehensive evaluation.
+The results for this method of training also provides a baseline against which to judge the effectiveness of other methods that have been proposed to improve generalisation performance.
+- noise
+Augmenting training images with random noise has been shown to reduce over-fitting, and hence, improve generalisation (Sietsma and Dow,, 1991; Holmstrom and Koistinen,, 1992; Goodfellow et al.,, 2016), especially to common image corruptions (Rusak et al.,, 2020; Gilmer et al.,, 2019).
+Various methods of adding noise to the training data have been proposed.
+Rusak et al., (2020) augmented half the training data with Gaussian noise, and found that this technique produced performance on corrupt test data that was competitive with the more complex augmentation method, AugMix, and that for the MNIST data-set specifically was just as effective as their proposed method of Adversarial Noise Training.
+Lopes et al., (2019) advocated adding Gaussian noise to only a randomly chosen patch of the image, which they claimed produced better robustness to image corruptions on CIFAR10 data than applying random noise to the whole image.
+Gilmer et al., (2019) used an alternative method for corrupting training images with Gaussian noise in which the standard deviation of the noise applied to each image was chosen with equal probability from the range .
+The method implemented here is a generalisation for the method proposed by Gilmer et al., (2019).
+Specifically, for each image the standard deviation of the noise was chosen from a uniform distribution with range .
+Then each pixel intensity value was independently modified by the addition of a value chosen at random from a Gaussian distribution with mean of zero and the chosen standard deviation.
+The original images had pixel intensity values in the range [0,1], and after the application of the noise pixel values were clipped at 0 and 1 to remain in this range.
+Appropriate values of and were found via a grid search using small ConvNets.
+Using this method and were chosen for use with CIFAR10, and the same hyper-parameters were also used for CIFAR100 and TinyIN.
+For MNIST was used.
+- AT
+Multi-step adversarial training with Projected Gradient Descent (PGD; Madry et al.,, 2018) is a standard, and highly effective, defence against adversarial attack.
+As a result, this method has become a standard benchmark against which all other methods of adversarial defence are judged.
+This method was, therefore, also selected as a data augmentation technique to be tested.
+It was implemented using the torchattacks PyTorch library (Kim,, 2021).
+To perform adversarial training, clean images from the current batch of training examples were used to generate non-targeted adversarial examples, that then replaced the original training images (as in Madry et al.,, 2018).
+Following standard protocols, the maximum allowed perturbation was constrained by the -norm to be less than the perturbation budget (), and the training-time adversarial images were generated using 10 steps of PGD (denoted as PGD for brevity).
+Standard values used in the previous literature were used for .
+Specifically, was set to for performing AT with the CIFAR10, CIFAR100 and TinyIN training data, and was used for MNIST.
+- pixmix
+This is a training data augmentation pipeline that repeatedly mixes (either additively or multiplicatively) an augmented image with an augmented clean image or a fractal image (Hendrycks et al.,, 2022).
+It is claimed that this augmentation method can improve generalisation to both clean and corrupt data, improve adversarial robustness, and improve unknown class rejection.
+This method of data augmentation was therefore also tested using the proposed, comprehensive, evaluation method.
+It was implemented using the official code released by the original authors.
+- regmixup
+A.2 Training Setup
+The ResNet18 (He et al.,, 2016) architecture was chosen as this is very commonly used in the literature.
+This model has just over 11M trainable parameters.
+Training was performed using a batch size of 128 for 110 epochs using the SGD optimiser with weight decay of 5e-4 and an initial learning rate of 0.1.
+The learning rate was step-wise decayed by a factor of 10 at the start of epochs 100 and 105.
+The set-up was based on that recommended by Pang et al., (2021) for performing AT on the CIFAR data-sets.
+Hence, these hyper-parameters may not be optimal for training with the other data augmentations used.
+For each experimental condition (combination of data-set, network architecture, and training data augmentation method) the network was trained and evaluated three times, each time with a different random weight initialisation and random presentation order of training samples.
+The reported result for each condition is the mean accuracy over these three trials.
+The standard-deviation in the mean performance measured using the proposed metric is shown in the tables of results and generally shows a high degree of consistency across trials.
+Publicly available checkpoints created by other researchers were only evaluated once.
+Open-source code written in Python using the Pytorch library (Paszke et al.,, 2019), which can be used to train networks and perform all the experiments described in this article, is available for download from: https://codeberg.org/mwspratling/RobustnessEvaluation.